### v1

# SCHEME: Scalable Channer Mixer for Vision Transformers

Abstract
Vision Transformers have received significant attention due to their impressive performance in many vision tasks. While the token mixer or attention block has been studied in great detail, the channel mixer or feature mixing block (FFN or MLP) has not been explored in depth albeit it accounts for a bulk of the parameters and computation in a model. In this work, we study whether sparse feature mixing can replace the dense connections and confirm this with a block diagonal MLP structure that improves the accuracy by supporting larger expansion ratios. To improve the feature clusters formed by this structure and thereby further improve the accuracy, a lightweight, parameter-free, channel covariance attention (CCA) mechanism is introduced as a parallel branch during training. This design of CCA enables gradual feature mixing across channel groups during training whose contribution decays to zero as the training progresses to convergence. This allows the CCA block to be discarded during inference, thus enabling enhanced performance with no additional computational cost. The resulting Scalable CHannEl MixEr (SCHEME) can be plugged into any ViT architecture to obtain a gamut of models with different trade-offs between complexity and performance by controlling the block diagonal structure size in the MLP. This is shown by the introduction of a new family of SCHEMEformer models. Experiments on image classification, object detection, and semantic segmentation, with different ViT backbones, consistently demonstrate substantial accuracy gains over existing designs, especially under lower FLOPs regimes. For example, the SCHEMEformer establishes a new SOTA of 79.7% accuracy for ViTs using pure attention mixers on ImageNet-1K at 1.77G FLOPs.
1 Introduction
Figure 1: Left: Accuray-FLOPs tradeoff for various SCHEME models (denoted in blue) based on popular ViT architectures. Right: Accuracy-FLOPs trade-off of the proposed SCHEMEformer family derived from the Metaformer-PPAA-S12 model [57] with higher expansion ratios in the MLP blocks. The SCHEMEFormer family outperforms the baseline for the same FLOPs, with larger gains for smaller models.
[…]
Figure 2: Proposed SCHEME channel mixer. The channel mixer of the standard transformer consists of two MLP layers, performing dimensionality expansion and reduction by a factor of $E$. SCHEME uses a combination of a block diagonal MLP (BD-MLP), which reduces the complexity of the MLP layers by using block diagonal weights, and a channel covariance attention (CCA) mechanism that enables communication across feature groups through feature-based attention. This, however, is only needed for training. The weights $1-\alpha$ decay to zero upon training convergence and CCA can be discarded during inference, as shown on the right. Experiments show that CCA helps learn better feature clusters, but is not needed once these are formed.
Vision Transformers (ViTs) [15, 33, 51, 32] have become ubiquitous in computer vision in recent years. Under this architecture, an image is decomposed into a set of patches which are considered as tokens and fed to a transformer model [49] of two main components: a spatial attention module, which reweighs each token according to its similarity to the other tokens extracted from the image, enabling information fusion across large spatial distances, and a channel mixer that combines the feature channels extracted from all patches using a multi-layer perceptron (MLP or FFN). A bottleneck of this model is the quadratic complexity of the attention mechanism on the number of patches. Numerous ViT variants have been proposed in recent years to improve complexity and/or accuracy. These can be grouped into 1) improved attention mechanisms, and 2) hybrid architectures that replace attention or combine it with convolutions.
Much less research has been devoted to the channel mixer. Most models simply adopt the MLP block of [49], consisting of two layers, where channels are first expanded by a specified expansion ratio and then compressed to the original dimension. This is somewhat surprising since the mixer is critical for good transformer performance. For example, Dong et. al.  [14] showed that pure attention, without MLPs or residual connections, collapses doubly exponentially to a rank one matrix. Yu et. al. [57] also showed that without residual connections or MLP, the model cannot converge. DaViT [12] showed that replacing MLPs with more attention blocks (both spatial and channel attention) of equivalent computational complexity lowers the transformer accuracy. These observations show that the channel mixer is an indispensable component of ViTs.
In this work, we investigate the channel mixer module, both to quantify its contribution to ViT performance and to improve the complexity vs. accuracy trade-off of the transformer model. We start by studying the impact of the channel expansion performed in the mixer, by varying its expansion ratio. Transformer performance is shown to increase until it saturates for an expansion ratio beyond 8. However, because mixer (MLP) complexity is linear on the dimension of the intermediate representation, naive channel scaling with expansion ratios larger than 4 causes an explosion of parameters and computation, leading to models of large complexity and prone to overfit.
To achieve a better trade-off between dimensionality and computation, we seek to leverage a hardware efficient mixer with block diagonal structure [1, 10]. This consists of breaking the input and output feature vectors of a network layer into disjoint groups and performing matrix multiplications only within each group, as illustrated in Figure 2. We denote the resulting MLP block as Block Diagonal MLP (BD-MLP). Despite the lack of feature mixing across groups, we find that a transformer model equipped with the BD-MLP and a larger expansion ratio (to match both the parameter count and computation) achieves comparable or slightly higher accuracy than a baseline with the dense MLP. This suggests that the lower accuracy of the sparse block diagonal operations is offset by the gains of larger expansion ratios. Further analyzing the features learned by the different groups of the BD-MLP, we observe that they form feature clusters that are similar in terms of their discriminative ability to identify the target classes.
To learn better feature clusters, we seek a mechanism capable of restoring inter-group feature communication only during training without increasing parameters and computational overhead. For this, we propose to use the channel correlations among features. This is done by introducing a channel attention branch that computes the channel covariance matrix of the input features of the BD-MLP and is used to re-weigh its output features, as illustrated in Figure 2. This attention mechanism is denoted as the channel covariance attention (CCA) block. The reweighted features are then fused with the BD-MLP output by means of a weighted residual addition with learned weights $(\alpha,1-\alpha)$. The overall design, combining the sparse block diagonal structure of the BD-MLP, and the parameter-free CCA attention module, is extremely efficient in terms of parameters. Experiments also show that the fusion weight learned in the CCA branch, denoted as $1-\alpha$ in Fig. 2, gradually decays to zero during training. This happens consistently across all layers of the model and across model architectures. While the CCA is important for the formation of good feature clusters during training, it can be removed at inference without any loss, as illustrated in Figure 2. As a result, the model accuracy improves over simply using the BD-MLP mixer, but the computational complexity of model inference does not. This leads to an extremely efficient inference setup, both in terms of parameters and FLOPs. We empirically verify the hypothesis through ablations that study the evolution of attention branch weights over training and the class separability of the learned feature groups.
Since the combination of BD-MLP and CCA enables the design of models with larger expansion ratios, we denote the resulting mixer as the Scalable CHannEl MixEr (SCHEME). SCHEME is a generic channel mixer that can be plugged into existing ViT variants to realize larger expansion ratios and obtain effective scaled-down or scaled-up model versions. We replace the channel mixer of the MetaFormer-PPAA-S12 [57] model with SCHEME to obtain a new family of SCHEMEformer models with improved accuracy/complexity trade-off. This is shown in the right of Figure 1. For the same parameter count and FLOPs, SCHEME obtains better accuracy than the baseline by using larger MLP expansion ratios. The gains are particularly larger for the smaller models. For example, the SCHEMEformer family achieves SOTA results of 79.74% accuracy at 1.77 G FLOPs, for ViTs using self-attention mixers. The left of Figure 1 shows that SCHEME models based on popular ViT architectures have a better Accuracy-FLOPs trade-off than the baseline models. Most of the SCHEMEformer models lie on the Pareto frontier of accuracy vs complexity, showing that SCHEME allows fine control over the two variables, while guaranteeing SOTA performance. These properties are shown to extend to object detection and semantic segmentation tasks, as well as to different architectures, such as T2T-ViT [60], CoAtNet [9], Swin Transformer [32], CSWin [13] and DaViT [12]. In summary, our contributions are as follows,
  • We study the channel mixer of MLPs in ViT and show that the dense feature mixing can be replaced by sparse feature mixing with higher internal feature dimensionality for improved accuracy.
  • We propose the SCHEME module, combining 1) a BD-MLP to enable internal feature representations of larger dimensionality than previous MLP blocks, and 2) CCA to enable the learning of these representations without cost at inference.
  • SCHEME is shown to be generic and easily plugged into existing ViTs to obtain a gamut of models with improved performance. We use it to introduce a new family of SCHEMEformer models with better accuracy/complexity trade-off than previous ViTs.
  • Experiments on image classification, object detection, and semantic segmentation consistently show gains in accuracy for fixed computation and model size.
2 Related Work
Vision Transformers: Vision transformers advanced the state-of-the-art in several vision tasks since [33, 15] successfully applied the transformer-based self-attention NLP model of [49] to image generation and classification tasks. These models rely on a spatial attention mechanism, based on the matrix of dot-products between features extracted from image patches. This has quadratic complexity in the number of patches and is quite intensive. Most follow up work [19, 40, 59, 45, 41, 54, 43, 60, 62, 50] improved the spatial attention mechanism of ViT. DeiT [43] and subsequent works [46, 47] introduced a distillation token to distill information, typically from a CNN teacher, without large amounts of data or compute. PvT [51] proposed a progressive shrinking pyramid architecture with spatial-reduction attention that scales ViTs for dense prediction tasks beyond image classification. Swin transformers [32] introduced a hierarchical shifted window attention mechanism, which reduces the complexity to linear with respect to the number of windows. HaloNet [50] proposed two extensions for local ViTs [34] with blocked local attention and relaxed translational equivariance for scaling ViTs. A more extensive review of ViTs is given in [31].
Recently, several works have shown that spatial attention is not the critical ViT feature. Some works improved performance by relying on hybrid architectures, which augment or replace ViT layers with convolutions [53, 48]. Efficient transformer designs, for edge devices, frequently sacrifice spatial attention to achieve better trade-offs between FLOPs and accuracy [5, 27]. Other works have questioned the need for spatial attention altogether. Metaformer [57] argued that the fundamental trait of ViT is the mixing of information across patches, showing that competitive results can be obtained by simply replacing attention with pooling or identity operations. Similarly, [30] showed that spatial attention is not critical for vision transformers by proposing a spatial-gating MLP of comparable performance to ViT [15]. DaViT [12] showed that it is helpful by simply reusing the spatial attention design for channel attention, building a cascade of alternating spatial and channel attention blocks. While this improves performance, it increases the complexity of the transformer block, resulting in a model with many parameters and potentially redundant channel mixer operations. Despite all this work on ViT architectures, little emphasis has been devoted to the other main component of the transformer: the channel mixer module (MLP) that follows attention. This is surprising because the mixer dominates both the parameter count and complexity (FLOPs) of the standard transformer block. [26, 59] modify the MLP block to mimic the inverted residual block of the MobileNetV2 [36], by adding a depthwise convolution. This improved performance but increases parameter and computation costs. Switch Transformer [17] replaces the FFN with sparse mixture of experts [39] that dynamically routes the input tokens. This design allows scaling models to large sizes using higher number of experts, however its not effective for ViTs. In this work, we instead propose an efficient and generic channel mixer module that improves both the parameter and computational efficiency of the transformer and allows for flexible scaling of ViTs.
3 The SCHEME module
Figure 2 depicts the proposed SCHEME module for feature mixing in ViTs. As shown in the figure, the standard channel mixer consists of two MLP layers, which expand the dimensionality of the input features and then reduce it to the original size. Let $\mathbf{x}\in\mathbb{R}^{d\times N}$ be the matrix containing the $N$ $d$-dimensional input feature vectors extracted from $N$ image patches. The mixer computes an intermediate representation $\mathbf{z}\in\mathbb{R}^{Ed\times N}$ and an output representation $\mathbf{y}\in\mathbb{R}^{d\times N}$ according to
$\mathbf{z}=\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1}{\bf 1}^{T}_{N})$ (1)
$\mathbf{y}=\mathbf{W}_{2}\mathbf{z}+\mathbf{b}_{2}{\bf 1}^{T}_{N}$ (2)
where $\mathbf{W}_{1}\in\mathbb{R}^{Ed\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times Ed},\mathbf{b}_{1}\in\mathbb{R}^{d},\mathbf{b}_{2}\in\mathbb{R}^{Ed},$ ${\bf 1}_{N}$ is the $N$-dimensional vector containing ones as all entries, $\sigma(.)$ is the activation function (typically GELU), and $E$ is an expansion factor, typically 4.
Table 1 details the impact of the mixer on overall transformer performance by evaluating the role of the expansion factor $E$ on the performance of MetaFormer-PPAA-S12 [57] (Pooling, Pooling, Attention, Attention) architecture on ImageNet-1K. Classification accuracy increases from 76.0 to $81.8\%$, as $E$ ranges from $1$ to $6$, decreasing for $E=8$, which suggests overfitting. The table also shows that these gains are not trivial. The MetaFormer-S18, which has no expansion ($E=1$) but more transformer layers and complexity comparable to that of the MetaFormer-S12 ($E=4$), has a significantly lower accuracy of 78.3%. In summary, for a given computation budget, it is beneficial to trade off transformer depth for dimensionality expansion in the channel mixer. This shows that this expansion is a critical component of the transformer architecture. On the other hand, naively scaling $E$ beyond 6 severely increases parameters and computation, leading to models that over-fit and are impractical for many applications.
3.1 Scalable Channel Mixer (SCHEME)
To efficiently realize the gains of higher intermediate MLP feature dimensionality, we propose the SCHEME module, which consists of a BD-MLP and CCA block where the CCA branch is only used during training. The following sections describe these in detail.
Block Diagonal MLP (BD-MLP): Block diagonal matrices have been previously proposed to efficiently approximate dense matrices [1, 10]. In CNNs, group channel operations are frequently used to design lightweight mobile models with improved accuracy-computation trade-off [25, 6, 36]. This consists of splitting the feature vectors of (1)-(2) into disjoint groups, e.g. $\mathbf{x}$ into a set of $g_{1}$ disjoint features $\{\mathbf{x}_{k}\}_{k=1}^{g_{1}}$ where $\mathbf{x}_{k}\in\mathbb{R}^{d/g_{1}\times N}$, and $\mathbf{y}$ into a set $\{\mathbf{y}_{k}\}_{k=1}^{g_{2}}$ where $\mathbf{y}_{k}\in\mathbb{R}^{Ed/g_{2}\times N}$. As illustrated in Figure 2, the MLPs of (1)-(2) are then implemented independently for each group, according to
$\mathbf{z}_{k}=\sigma(\mathbf{W}_{1,k}\mathbf{x}_{k}+\mathbf{b}_{1,k}{\bf 1}^{T}_{N})$ (3)
$\mathbf{y}_{k}=\mathbf{W}_{2.k}\mathbf{z}_{k}+\mathbf{b}_{2,k}{\bf 1}^{T}_{N}$ (4)
where $\mathbf{W}_{1}\in\mathbb{R}^{Ed/g_{1}\times d/g_{1}},\mathbf{W}_{2}\in\mathbb{R}^{d/g_{2}\times Ed/g_{2}},\mathbf{b}_{1}\in\mathbb{R}^{d/g_{1}},\mathbf{b}_{2}\in\mathbb{R}^{Ed/g_{2}}$ and $\mathbf{z}$ is decomposed into a set $\{\mathbf{z}_{k}\}_{k=1}^{G}$ where $\mathbf{z}_{k}\in\mathbb{R}^{d/G\times N}$, with $G=g_{1}$ in (3) and $G=g_{2}$ in (4). Since the complexity of (3) is $g_{1}^{2}$ times smaller than that of (1) and there are $g_{1}$ groups, the complexity of the first MLP is $1/g_{1}$ times that of standard MLP. Similarly, the complexity of the second MLP is $1/g_{2}$ times that of the standard MLP. Hence, a transformer equipped with the BD-MLP and expansion factor $\frac{2g_{1}g_{2}}{g_{1}+g_{2}}E$ has identical complexity to a standard transformer of factor $E$. For example, when $g_{1}=g_{2}=g$ this allows growing the expansion factor by a factor of $g$ without computational increase.
Expansion FLOPs Top-1 Acc
Model ratio ($E$) (G) (%)
[…]
Metaformer-PPAA-S12 [57] 4 2.55 81.0 (+0.0)
[…]
Table 1: Effect of the MLP block expansion ratio ($E$) on ImageNet-1K validation accuracy for the MetaFormer-S12 [57] architecture. *{*}* denotes results obtained with the code made publicly available by the authors.
[…]
$CCA(\mathbf{x})=\mbox{softmax}\left(\frac{\mathbf{x}\mathbf{x}^{T}}{\tau}\right)\mathbf{x}$ (5)
where the softmax operation is applied across the matrix rows and $\tau$ is a smoothing factor. The output of the channel mixer block is obtained by the weighted summation of the BD-MLP and CCA branches, according to
$\begin{split}\textbf{y}_{out}=\alpha\textbf{y}+(1-\alpha)CCA(\textbf{x})\end{split}$ (6)
where $\alpha$ is a mixing weight learned across all samples. Various other design choices are discussed in Section 4.2.
CCA as a Regularizer: The design choice of using a parameter free attention branch and a static weight $\alpha$ allows the model to form better feature clusters during training and gradually decay the contribution from CCA branch once the feature clusters are formed. This can be seen in Figure 3(a), which plots the value of the learned mixing weight $1-\alpha$ as a function of training epochs on ImageNet-1K, for all transformer layers. These plots are typical of the behavior we observed with all transformer backbones and architectures we considered. Clearly, $1-\alpha$ starts with high to intermediate values, indicating that information flows through both branches of the mixer, but decays to $1-\alpha\approx 0$ as training converges. Hence, there is no degradation if the CCA branch is removed during inference. This eliminates a substantial amount of computation during inference, leading to the the training and inference setup of Figure 2, where CCA is not used at inference.
We explain this behavior by conjecturing that the downside of the computational efficiency of the BD-MLP is a more difficult learning problem, due to the independent processing of channel groups. This creates symmetries in the cost function, e.g. the order of the feature groups is not important, and requires a feature clustering operation that is likely to produce more local minima. The CCA branch helps to smooth out this cost function during training, while the feature groups are not established, by allowing inter-group communication. However, once the right feature groupings are found, CCA is no longer needed, and a simple BD-MLP mixer has no loss of performance over the standard MLP. Note that the operation of (5) is basically a projection of $\mathbf{x}$ into canonical subspaces of features that are correlated in the input image. This is likely to be informative to guide the group formation, but less useful when the features are already clustered. While this hypothesis is not trivial to test, since (5) varies from example to example, we confirmed that using CCA during training enhances class separability, which likely reduces overfitting for large expansion ratios. See the results section for more details.
3.2 The SCHEMEformer family
The proposed SCHEME module can be plugged into any ViT architecture, enabling efficient control of model complexity via the mixer hyperparameters $g_{1}$, $g_{2}$, and $E$. Table 1 shows that naively scaling down the ViT model by simply reducing $E$ causes a significant accuracy loss. The SCHEME module allows much more effective control of the accuracy/complexity trade-off, producing models of better performance for a fixed computational budget. This is demonstrated by the introduction of a new family of models, denoted as SCHEMEformer, obtained by replacing the channel mixer of the Metaformer-PPAA [57] architecture with the SCHEME module. Two such configurations are shown in Table 2 where the naming follows the convention {model-name}-{$g_{1}$$g_{2}$}-e{$E$}. These are shown to achieve SOTA accuracy for lower computational budgets in the experiments section, where we also discuss results with other backbones.
S12 S24 S36
P F P F P F
Model (M) (G) (M) (G) (M) (G)
SCHEMEformer-PPAA-44-e8 12 1.8 21 3.3 55 7.96
SCHEMEformer-PPAA-12-e8 21 3.3 40 6.5 59 9.6
Table 2: Scaling MetaFormer-PPAA using SCHEME. Each column of the table reports to one baseline model (S12 to S36) of increasing size. P and F denote Parameters and FLOPs respectively. SCHEMEformer-PPAA-44-e8 (SCHEMEformer-PPAA-12-e8) models use configurations $g_{1}$, $g_{2}$, and $E$ that downscale (upscale) the baseline.
3.3 Computational Complexity
The complexity of the BD-MLP block is controlled by the group numbers $g_{1}$, $g_{2}$ and the expansion factor $E$, with a total cost $\mathcal{O}(Ed^{2}/g_{1}+Ed^{2}/g_{2})$, where $d$ is the channel dimension. The computational cost of CCA is $\mathcal{O}(Nd^{2})$ where $N$ is the number of tokens. Since CCA is not used during inference, it only adds to the computations during training. The SCHEME framework provides a systematic way to control the trade-off of transformer width vs depth, by controlling the block size and expansion hyperparameters.
4 Experimental Results
In this section, we discuss an evaluation of the SCHEME module on Image Classification (ImageNet-1K [11]), Object Detection (COCO-17 [28]), and Semantic Segmentation (ADE-20K [65]) benchmarks.
#Params FLOPs Top-1 Acc
Model  (M) (G) (%)
Attention Mixers
Poolformer-S7 [57] 8.60 1.1 73.0
PVT-Tiny [51] 13 1.9 75.1
XCiT-T12 [16] 7 1.2 77.1
Poolformer-S12 [57] 12 1.8 77.2
MetaFormer-PPAA-11-e2-S12 [57] 12 1.8 78.9
XCiT-T24 [16] 12 2.3 79.4
ViT-B/16 [15] 86 17.6 79.7
SCHEME-Poolformer-44-e8-S12 7.20 1.0 73.0
SCHEME-Poolformer-12-e8-S12 16.7 2.6 78.5
SCHEME-CSWin-44-e8-Tiny 15.6 3.0 79.2
SCHEME-Swin-44-e8-Tiny 19.7 3.1 79.6
SCHEMEformer-PPAA-44-e8-S12 12 1.77 79.7
PVT-Small [51] 25 3.8 79.8
QuadTree-B-b1 [41] 14 2.3 80.0
Poolformer-S24 [57] 21.0 3.4 80.3
MetaFormer-PPAA-S12 [57] 17 2.6 81.0
PVT-Medium [51] 44 6.7 81.2
Swin-Tiny [32] 29 4.5 81.3
T2T-ViT t-14 [60] 22 6.1 81.7
DeiT-B [43] 86 17.5 81.8
CaiT-XS-24 [45] 27 5.4 81.8
ViL-Small [61] 25 5.1 82.0
XCiT-S12 [16] 26 4.8 82.0
SCHEME-Poolformer-12-e8-S24 30.8 4.9 80.5
SCHEME-T2T-ViT-44-e8-14 17.4 6.1 81.7
SCHEME-Swin-12-e8-Tiny 36.9 5.9 81.7
SCHEMEformer-PPAA-44-e8-S24 21 3.31 81.9
SCHEMEformer-PPAA-12-e8-S12 21 3.35 82.0
SCHEME-T2T-ViT-12-e8-14 27.7 8.1 82.1
SCHEME-DaViT-44-e8-Tiny 19.7 3.8 82.1
Focal-Tiny [56] 29 4.9 82.2
CrossViT-Base [2] 105 21.2 82.2
T2T-ViT-24[60] 64 14.1 82.3
CPVT-Base [7] 88 17.6 82.3
DaViT-Tiny [12] 23.0 4.3 82.8
CSWin-Tiny [13] 23.0 4.3 82.8
SCHEMEformer-PPAA-12-e8-S24 40 6.47 82.8
SCHEME-DaViT-12-e8-Tiny 37.0 6.6 83.0
SCHEME-CSWin-12-e8-Tiny 29.1 5.6 83.2
XCiT-L24 [16] 189 36.1 82.9
TNT-B [20] 66 14.1 82.9
Swin-Small [32] 50 8.7 83.0
ViL-Base [61] 56 13.4 83.2
CSWin-Small [13] 35.0 6.9 83.6
PVTv2-B5 [52] 82.0 11.8 83.8
Focal-Base [56] 89.8 16.0 83.8
SCHEMEformer-PPAA-44-e8-S36 55 7.96 83.2
SCHEMEformer-PPAA-12-e8-S36 58 9.58 84.0
Convolution+Attention Mixers
CeiT-S [59] 24 4.8 82.0
CvT-13 [53] 20 4.5 81.6
CoAtNet-0 [9] 25 4.2 81.6
PatchConv-S60 [44] 25 4.0 82.1
CvT-21 [53] 32 7.1 82.5
ViTAEv2-S [63] 19 5.7 82.6
$*{}^{*}$ 17.8 3.4 80.7
$*{}^{*}$ 24 4.1 81.6
SCHEME-CAformer-44-e8-S18 19 2.92 82.4
SCHEME-CAformer-12-e8-S12 24 3.61 82.9
Table 3: Image Classification on ImageNet-1K dataset. Comparison with SOTA ViT models grouped by the type of token mixers and accuracy. Proposed SCHEMEformer-PPAA-44-e8(12-e8) models are shown in light blue and the baseline architectures are shown in light cyan. *{*}* denotes our implementation as their code is not publicly available.
Figure 3: Comparison of various SCHEMEformer (blue stars) and SOTA ViT architectures (orange circles) (calculated on NVIDIA-Titan-X GPU with batch size 1). Left: Accuracy-Throughput trade-off. Right: Accuracy-Parameters trade-off.
4.1 Comparisons to the state of the art
Implementation Details: For image classification, we use a batch size of 1024, adjusted according to memory availability, and train all models on 8 GPUs with AdamW optimizer and cosine annealing learning rate schedule, of learning rate 1e-3. Warmup learning is applied in the first 5 epochs. Object detection is based on the setting of [57], where all backbones are trained with RetinaNet [29] and Mask-RCNN [22] detection heads, for a fair comparison with prior methods. A 1x learning schedule (12 epochs with AdamW optimizer and initial learning rate of 1e-4) is used and all models are implemented with the mmdetection [3] codebase, capping input sizes at 800, 1333 for the shorter and longer sides, respectively. For semantic segmentation, following prior methods [4, 24, 57], we train with a batch size of 16 for 80,000 iterations with AdamW optimizer and an initial learning rate of 2e-4. We use the mmsegmentation [8] codebase with an input size $512\times 512$ for training and resize the shorter side to 512 pixels for testing. For both detection and segmentation, imagenet pretrained models are used to initialize the backbones.
#Par $AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$ RetinaNet 1× ResNet-18[21] 21.3 31.8 49.6 33.6 16.3 34.3 43.2 PoolFormer-S12[57] 21.7 36.2 56.2 38.2 20.8 39.1 48.0 ResNet-50[21] 37.7 36.3 55.3 38.6 19.3 40.0 48.8 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 PoolFormer-S24[57] 31.1 38.9 59.7 41.3 23.3 42.1 51.8 ResNet-101[21] 56.7 38.5 57.8 41.2 21.4 42.6 51.1 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 Mask R-CNN 1× #Par $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ Mask R-CNN 1× ResNet-18[21] 31.2 34.0 54.0 36.7 31.2 51.0 32.7 PoolFormer-S12[57] 31.6 37.3 59.0 40.1 34.6 55.8 36.9 ResNet-50[21] 44.2 38.0 58.6 41.4 34.4 55.1 36.7 SCHEMEformer-44-e8-S12 31 39.8 61.9 42.9 24.1 53.0 42.3 PoolFormer-S24[57] 41.0 40.1 62.2 43.4 37.0 59.1 39.6 ResNet-101[21] 63.2 40.4 61.1 44.2 36.4 57.7 38.8 SCHEMEformer-44-e8-S24 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 4: COCO-17 Object Detection and Instance Segmentation. All backbones are pretrained on ImageNet-1K and use 1x learning schedule. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
Backbone #Par FLOPs mIoU (%) ResNet-18[21] 16 32.2 32.9 PVT-Tiny[51] 17 33.2 35.7 ResNet-50[21] 29 45.6 36.7 PoolFormer-S12[57] 16 30.9 37.2 ResNet-101[21] 48 65.1 38.8 ResNeXt-101-32x4d[55] 47 64.7 39.7 PVT-Small[51] 28 44.5 39.8 XCiT-T12/8[16] 8.4 - 39.9 ResNeXt-101-64x4d[55] 86 103.9 40.2 PoolFormer-S24[57] 23 39.3 40.3 SCHEMEformer-44-e8-S12 15.5 34.3 40.9 PVT-Medium[51] 48 61.0 41.6 PoolFormer-S36[57] 35 47.5 42.0 PVT-Large[51] 65 79.6 42.1 PoolFormer-M36[57] 60 67.6 42.4 SCHEMEformer-44-e8-S24 24.8 45.7 42.5 Table 5: Semantic Segmentation results on ADE20K dataset using Semantic FPN [24]. All backbones are pretrained on ImageNet-1K. FLOPs are calculated at $512\times 512$ resolution.
Model BDM CCA Acc (%) Baseline 78.9 44-e8-S12 79.1 (+0.2) 44-e8-S12 79.7 (+0.6) Table 6: Ablation study of the contribution of BD-MLP and CCA branches to the overall performance for SCHEMEformer-PPAA-44-e8-S12 model. Baseline refers to Metaformer-PPAA-11-e2-S12. Module #Par FLOPs Top-1  (M) (G) Acc (%) Shuffle 11.83 1.77 79.1 SE 11.98 1.77 79.3 Conv 13.31 1.97 79.6 DyCCA 11.90 1.83 79.6 CCA 11.83 1.77 79.7 Table 7: Ablation study of alternative designs for feature attention branch in SCHEME.
Image Classification: Image classification is evaluated on Imagenet-1K, without using extra data. We report the results with single crop top-1 accuracy at $224\times 224$ input resolution. Results are presented for various backbones and hyperparameter configurations, to show the generalizability of the SCHEME module. We evaluate the SCHEMEformer family of models based on the Metaformer-PPAA-S12 [57] obtained by replacing the MLP of the original model with the SCHEME module. Additionally, we replace the pooling of the first two stages of the Metaformer-PPAA-S12 with convolution  [58] to evaluate SCHEME-CAformer models.
Table 3 compares SCHEME models to various SOTA models from the literature grouped by the type of token mixers and accuracy. The SCHEME models are highlighted in blue and the baseline models on which they are based in cyan. While the table is ordered by increasing accuracy, meaningful comparisons must account for both accuracy and parameter size or FLOPs, since the models have very different sizes. The SCHEMEformer family achieves the best trade-off between accuracy and computation or parameter size by using a larger internal feature dimension for the MLP where the gains are larger for smaller models, whose performance is not near saturation. For example, as highlighted in red font, for the same compute and parameter cost, the SCHEMEformer-PPAA-44-e8-S12 improves on the performance of the baseline MetaFormer-PPAA-11-e2-S12 by +0.8% and achieves a SOTA result of 79.7% top-1 accuracy at 1.77G FLOPs. The table also shows the accuracy of the SCHEME module applied to other popular ViT architectures, such as T2T-ViT, Swin, CSWin, Da-ViT in the Attention Mixers category, and CoatNet, CAformer in the Convolution+Attention Mixers category realizing models of intermediate complexity with comparable or better accuracy for the same parameters and FLOPs. For example, SCHEME-CoAtNet-12-e8-0 achieves 81.6% accuracy with 1M fewer parameters and 0.1G fewer FLOPs than the baseline CoAtNet-0 (highlighted in brown). SCHEME-CSWin-12-e8-Tiny (blue) achieves 83.2% accuracy with just 29 M parameters and 5.6 G FLOPs sandwiched between CSWin-Tiny (82.8%) and CSWin-Small (83.6%).
Figure 1 provides a visual illustration of the results in the table. The plot on the left compares the SCHEME family (blue stars) to all other models (orange dots) in terms of the trade-off between FLOPS and accuracy. It is clear that most SCHEME models lie on the Pareto frontier for these two objectives, significantly expanding the set of models on this frontier. This illustrates the fine control that SCHEME allows over the accuracy/complexity trade-off of transformer models. The plot on the right shows how the SCHEMEFormer models improve on the accuracy/complexity trade-off of the MetaFormer model on which they are based. This shows that the SCHEME module generalizes well to different backbones and offers higher flexibility in scaling.
Throughput and Memory: Beyond higher accuracy, SCHEMEformer models also have low latency and parameter counts. The left of Figure 3 shows a favorable throughput vs accuracy trade-off for various SCHEME models compared to SOTA ViT architectures. Throughput is computed on NVIDIA-Titan-X GPU with a batch size of 1 averaged over 1000 runs. Despite having larger internal dimensions, the SCHEMEformer-PPAA-12-e8-S12 has a throughput (130 images per second) comparable to that of the baseline MetaFormer-PPAA-11-e2-S12. This is a consequence of the block diagonal structure, which enables higher FLOPs utilization [18]. The right of Figure 3 shows a favourable accuracy vs parameters tradeoff comparison for SCHEME models compared to SOTA ViT architectures. SCHEME based scaling achieves higher accuracy with lower parameter counts. These results show that SCHEME models lie on the Pareto frontiers of accuracy vs throughput and vs memory consumption.
Object Detection: Table 4 compares SCHEMEformer models to models of similar complexity on the COCO-17 object detection and instance segmentation benchmark, for both Retinanet and Mask-RCNN detection heads. The table is ordered by increasing mAP ($AP$, $AP^{b}$) values. Again, the SCHEMEformer models outperform all other models of the same or smaller size. For example, SCHEMEformer-44-e8-S12 with RetinaNet head achieves 38.3% overall mAP with just 21M params, outperforming models of comparable size by more than 2% and achieving comparable performance to models with $2.5$ times their size. Similar gains are obtained for the Mask-RCNN based SCHEMEformer-44-e8-S12. These results show that the SCHEME module also advances object detection.
Semantic Segmentation: Table 5 compares SCHEMEformer-44-e8 S12 and S24 models using the semantic FPN framework [24] to other models of similar complexity on ADE20K semantic segmentation benchmark. As before, the SCHEMEformer models outperform all other models of the same or smaller size. For example, SCHEMEformer-44-e8-S12 achieves 40.9% mIoU, which is 5.2% higher than the model of most comparable size, PvT-Tiny. Similarly, the S24 model outperforms PoolFormer-M36 using only 41% of its parameters. In general, the gains obtained for semantic segmentation are higher than those of object detection, suggesting that the feature clusters formed by the block diagonal structure have larger benefits for dense prediction tasks.
4.2 Ablation Studies
In this section, we discuss ablations of the contribution of the BD-MLP and CCA branches, alternative designs for the channel mixer, and the regularizing effect of CCA.
Contribution of BD-MLP and CCA branch: Table 6 shows an ablation of the contribution of the BD MLP and CCA branches to transformer performance. Starting from the Metaformer-PPAA-11-e2-S12, with expansion ratio $E=2$ and dense MLP, we replace the channel mixer with that of the SCHEME to obtain SCHEMEformer-PPAA-44-e8-S12, which has $g_{1}=g_{2}=4$ and $E=8$, maintaining the number of parameters and FLOPs fixed. The SCHEME model with only BD-MLP improves on the baseline by 0.2%. The addition of the CCA branch provides an additional gain of 0.6%, showing the gains of better feature clusters. Since CCA is not used at inference, its gains are free in terms of additional parameters/FLOPs.
Effect of large expansion ratios: The right of Figure 1 shows the effect of simultaneously increasing the expansion ratio $E$ and adjusting groups to realize different models of similar size and complexity. All models are based on the Metaformer-PPAA-S12. For fixed parameters/FLOPs, SCHEMEformer models achieve a gain of 1.0% to 1.4% over the baseline by increasing $E$ from 1 to 4. The gains increase with larger expansions and saturate at larger FLOPs. This shows that higher internal feature dimensions are crucial for obtaining better accuracy with smaller ViT models.
Alternative designs of Channel Mixer: We investigate whether alternative choices to the CCA branch could accomplish this goal more effectively. Table 7 compares the performance of models that replace CCA with other feature mixing operations: the channel shuffling operation of ShuffleNet [64], a squeeze and excitation [23] network (SENet), a single layer of convolution, and a dynamic version of CCA (DyCCA), where the weight $\alpha$ of (6) is predicted dynamically, using GCT attention [35]. The table shows that CCA obtains the best result.
It is worth noting that while CCA has a lot more computation than some of these alternatives, it is not needed during inference. We found this not to be true for the alternatives, which produce much more balanced weights $\alpha$ after training convergence, and cannot be discarded at inference without a performance drop. We conjecture that, because all these alternatives have learnable parameters, the network learns to leverage them to extract complementary features, which must be fused with those produced by the BD-MLP branch during inference. Thus, while the alternative designs are very small modules, they incur an additional computational cost. CCA does not, achieving the best performance with no additional cost during inference.
[…]
Regularizing effect of CCA: Fig. 3(a) shows the evolution of the weight $1-\alpha$ of the CCA branch in (6) during training. While initially large, it gradually decays to zero as training progresses. The figure shows that this holds for all network layers. Hence, CCA can be discarded at inference. Fig. 5 in appendix plots the weights $1-\alpha$ upon training convergence, for the family of SCHEMEformer-PPAA-44-e8 models, confirming that the weights are indeed very close to zero across all layers. Table 8 shows the impact of removing CCA at inference in top-1 accuracy. The performance decrease is negligible ($\approx 0.02$) across backbones.
[…]
(a) Evolution of weight $1-\alpha$
(b) Class separability
Figure 4: Regularizing effect of CCA. Left: Evolution of weight $1-\alpha$ across all layers for SCHEMEformer-PPAA-44-e8-S12 model. Right: Comparison of class separability of the features $\bf y$ of the SCHEME module, for 50 random classes on ImageNet-1K validation set, using SCHEMEformer-PPAA-44-e8-S12 model trained with and without CCA.
We conjectured above that CCA helps training because it facilitates feature clustering into naturally independent groups that do not require inter-group communication. We tested this hypothesis by studying the intermediate feature vectors ${\bf v}_{l}$ (obtained after max-pooling the features $\bf y$ of (4)) extracted from four randomly selected layers $l$ of two ImageNet pretrained models, trained with and without CCA respectively. We split the features into the 4 groups used in the model ${\bf v}_{l,g},g\in\{1,\ldots 4\}$ and concatenated the features of all layers in the same group. This produced four vectors ${\bf u}_{g}=\mbox{concat}(\{v_{l,g}\}_{l})$ containing the features of each group $g$ extracted throughout the network. A linear classifier was then learned over each vector ${\bf u}_{g}$. Table 9 shows the top-1 accuracy per feature group and model. To evaluate whether groups learn different class clusters, we also average the outputs from the four group classifiers to obtain the final accuracy. Without CCA, i.e. no group communication during training, the network produces feature groups that tend to be individually more predictive of the image class, but less predictive when combined. This suggests that there is some amount of redundancy between the features extracted by the different groups. By introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
Fig. 3(b) shows the class separability of the intermediate features $\bf y$ of (4) of a randomly chosen layer of the SCHEMEformer-PPAA-44-e8-S12 model. Class separability was measured as in [37], with a final value obtained by averaging the class separability across all classes. The model trained with CCA has higher class separability than that without it. This confirms that CCA is helpful in forming feature clusters that increase class separability during training. Conversely, we tested if CCA is helpful when channel shuffling is inserted in between the two mixer MLP layers, which destroys the group structure. This variant of the SCHEMEformer-PPAA-44-e8-S12 model achieved an accuracy of 79.1% for both training with and without CCA. This shows that CCA is not helpful when feature groups are mixed. Similarly, CCA did not provide any gains when applied to the standard MLP branch with full feature mixing (one group). All these results suggest that CCA indeed helps to form the independent feature clusters needed to achieve the computational efficiency of channel groups without performance degradation.
5 Conclusion
In this work, we proposed the SCHEME module with sparse feature mixing for improving the performance of ViTs. This uses a weighted fusion of a BD-MLP branch, which abstracts existing MLPs with block diagonal structure, and a parameter-free CCA branch that helps to cluster features into groups during training. The CCA branch was shown to improve training but not be needed at inference. Experiments showed that it indeed improves the class separability of the internal feature representation of the BD-MLP branch, helping create feature clusters that are informative of the image class. The standard MLP in transformer block is replaced with the SCHEME module to obtain a new family of SCHEMEformer models that were shown to improve performance for classification, detection, and segmentation, for fixed parameters and FLOPs with favorable latency. SCHEME is shown to be effective for various ViT architectures and was also shown to provide a flexible way to scale models, always outperforming models with smaller MLP expansion ratios having the same complexity.
Acknowledgment
We acknowledge and thank the use of the Nautilus platform for some of the experiments discussed above.
Appendix
Appendix A Ablation Studies
CCA #Params FLOPs Top-1 Acc
Model Used  (M) (G) (%)
SCHEMEformer-PPAA-44-e8-S24 21.2 4.1 81.88
SCHEMEformer-PPAA-44-e8-S24 21.2 3.3 81.88
SCHEMEformer-PPAA-44-e8-S36 55.2 10 83.20
SCHEMEformer-PPAA-44-e8-S36 55.2 8.0 83.15
SCHEMEformer-PPAA-12-e8-S12 21.3 3.8 82.00
SCHEMEformer-PPAA-12-e8-S12 21.3 3.4 81.98
SCHEMEformer-PPAA-12-e8-S24 40.0 7.3 82.80
SCHEMEformer-PPAA-12-e8-S24 40.0 6.5 82.76
SCHEMEformer-PPAA-12-e8-S36 58.8 10.8 84.00
SCHEMEformer-PPAA-12-e8-S36 58.8 9.6 83.95
Table 10: Ablation studies on removing CCA branch during inference for the family of SCHEMEformer-PPAA models on ImageNet-1K dataset.
CCA Group1 Group2 Group3 Group4 Ensemble Model (%) (%) (%) (%) (%) SCHEMEformer-44-e8-S12 29.63 37.59 37.57 35.89 73.95 SCHEMEformer-44-e8-S12 27.89 37.64 30.02 37.67 75.04 Table 11: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module. We train a linear classifier on top of the four feature groups extracted from the final MLP mixer of the transformer block of the network. The model trained with CCA forms feature clusters that learn diverse and complementary set of features that can obtain 1% higher validation accuracy than the model trained without CCA.
[…]
Backbone CCA RetinaNet 1× Mask R-CNN 1× Used #P $AP$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP_{M}$ $AP_{L}$P $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 13: Ablation study of removing CCA for COCO-17 Object Detection and Instance Segmentation. Removing CCA at inference does not impact the AP values as they are identical to the model using CCA at inference. $AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
CCA Semantic FPN #Par FLOPs mIoU (%) 15.5 36.4 40.9 15.5 34.3 40.9 24.8 49.8 42.5 24.8 45.7 42.5 Table 14: Ablation study of removing CCA for Semantic Segmentation results on ADE20K dataset.
Model SCHEMEformer-PPAA-44-e8 SCHEMEformer-PPAA-12-e8 SCHEME-CAformer
S12 S24 S36 S12 S24 S36 44-e8-S18 12-e8-S12
Peak drop rate of stoch. depth $d_{r}$ 0.1 0.2 0.4 0.1 0.2 0.4 0.15 0.15
LayerScale initialization $\epsilon$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-5}$ $10^{-5}$ $10^{-6}$ $10^{-5}$ $10^{-5}$
Data augmentation AutoAugment
Repeated Augmentation off
Input resolution 224
Epochs 300
Hidden dropout 0
GELU dropout 0
Classification dropout 0
Random erasing prob 0.25
EMA decay 0
Cutmix $\alpha$ 1.0
Mixup $\alpha$ 0.8
Cutmix-Mixup switch prob 0.5
Label smoothing 0.1
Batch size used in the paper 1024
Peak learning rate used in the paper $4\times 10^{-4}$
Learning rate decay cosine
Weight decay 0.05
Gradient clipping None
Warmup epochs 5 20
Relation between peak learning rate and batch size lr = batch $\text{lr = }\frac{\text{batch size}}{1024}\times e^{-3}$ lr = batch $\text{lr = }\frac{\text{batch size}}{1024}\times 8\times e^{-3}$
Optimizer AdamW LAMB
Adam $\epsilon$ $1e^{-8}$ None
Adam ($\beta_{1}$, $\beta_{2}$) (0.9, 0.999) None
Table 15: Hyperparameter Settings for the family of SCHEMEformer models trained on ImageNet-1K dataset.
A.1 Plot of the learned $1-\alpha$ weights in the SCHEME module for SCHEMEformer models
(a) SCHEMEformer-44-e8-S12
(b) SCHEMEformer-44-e8-S24
(c) SCHEMEformer-44-e8-S36
(d) SCHEME-CoAtNet-44-e8-Tiny
(e) SCHEMEformer-12-e8-S12
(f) SCHEMEformer-12-e8-S24
(g) SCHEMEformer-12-e8-S36
(h) SCHEME-Swin-12-e8-Tiny
Figure 5: Plot of the learned weight (1-$\alpha$) values across different layers of a network for the family of SCHEMEformer models. The weights reach a peak value near the middle of the network. We demonstrate that these peak weights are not yet converged and training the network for more epochs decays these weights to zero while also improving the accuracy. For example, training the SCHEMEformer-PPAA-12-e8-S12 model for 200 additional epochs reduced the weight norm of the vector of $1-\alpha$ weights from 0.18 to 0.05 showing that these weights gradually approach zero as the training progresses while improving the accuracy further by 0.3%.
The learned weights $1-\alpha$ for the family of SCHEMEformer models is shown in Fig. 5. Interestingly, the learned weights coarsely approximate the shape of a gaussian distribution. The learned weights reach a peak value in the middle layers of the network and drop to zero for all the other layers. The middle of the network typically correspond to the initial few layers of the third stage of the model that contains the maximum number of transformer blocks for all the models shown in Fig. 5. We conjecture that the weights for these layers have not fully converged and that feature mixing can still be useful for these layers and so training for more epochs will allow the $1-\alpha$ weights of these layers to converge to zero. To test this hypothesis, we trained the SCHEMEformer-PPAA-12-e8-S12 model for an additional 200 epochs beyond the standard 300 epochs and observed that the peak value of the $1-\alpha$ weights decreased further by 0.11 as compared to the model trained for 300 epochs and the accuracy improved by 0.3%. The weight norm (of all layers) decreased from 0.18 for 300 epoch model to 0.05 for 500 epoch model. This confirms our hypothesis and training SCHEMEformer models for larger epochs can further improve the accuracy.
A.2 More results of removing CCA at inference
Table 10 shows the results of the full family of SCHEMEformer models with and without using CCA at inference. It is straightforward to observe that removing CCA does not impact the accuracy consistently across different model depths and larger models have slightly higher accuracy drop as compared to smaller models while still being negligible overall ($\leq 0.05\%$).
Table 13 and 14 show the results of removing CCA at inference for object detection and semantic segmentation models, respectively. For both tasks, results are identical to the model using CCA showing that the result also generalizes to downstream tasks.
A.3 CCA
In the paper, feature groups across different layers of the model was used to demonstrate the learning of feature clusters by CCA. Here, the feature from the final layer of the model is only considered. Table 11 shows the top-1 accuracy per feature group and model. The average of the outputs from the four group classifiers is reported in the final column of the table. The effect is more pronounced when using a single layer feature with more than 1% accuracy difference between the model with and without CCA. This further reinforces that by introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
A.4 Larger expansion ratios
Table 12 shows additional results of using larger expansion ratios with SCHEME for the same parameters and FLOPs using the Metaformer-PPAA-S12 baseline model. We observe that SCHEMEformer-PPAA consistently outperforms the baseline for larger expansions ratios and saturates as the model size and FLOPs increases.
Appendix B Implementation Details
B.1 Hyperparameter Settings
Table 15 shows the detailed hyperparameter settings of the family of SCHEMEformer models reported in the main paper.
Appendix C Qualitative Analysis: Grad-CAM [38] Visualization
Fig. 6 shows the results of class activation maps for SCHEMEformer-PPAA-44-e8-S12 model for few examples from the validation set of ImageNet-1K dataset. The stronger heatmap responses around the salient features of an object (e.g., body of a bird, cat) shows that the model ignores the background and attends to more discriminative spatial regions. Fig. 6 also shows the comparison with a few existing methods that demonstrates that SCHEMEformer-PPAA-44-e8-S12 attends to the complete object class and is better than the competing methods.
(a) Input Images
(b) ResNet-50 [21]
(c) ResMLP [42]
(d) DeiT-Small[43]
(e) Poolformer-S12[57]
(f) SCHEMEformer-PPAA-44-e8-S12
Figure 6: GRAD-CAM [38] visualization for a few validation samples on ImageNet-1K dataset for SCHEMEformer-PPAA-44-e8-S12 model.

### v4

@@
-SCHEME: Scalable Channer Mixer for Vision Transformers
+SCHEME: Scalable Channel Mixer for Vision Transformers
 Abstract
-Vision Transformers have received significant attention due to their impressive performance in many vision tasks.
-While the token mixer or attention block has been studied in great detail, the channel mixer or feature mixing block (FFN or MLP) has not been explored in depth albeit it accounts for a bulk of the parameters and computation in a model.
-In this work, we study whether sparse feature mixing can replace the dense connections and confirm this with a block diagonal MLP structure that improves the accuracy by supporting larger expansion ratios.
-To improve the feature clusters formed by this structure and thereby further improve the accuracy, a lightweight, parameter-free, channel covariance attention (CCA) mechanism is introduced as a parallel branch during training.
-This design of CCA enables gradual feature mixing across channel groups during training whose contribution decays to zero as the training progresses to convergence.
-This allows the CCA block to be discarded during inference, thus enabling enhanced performance with no additional computational cost.
-The resulting Scalable CHannEl MixEr (SCHEME) can be plugged into any ViT architecture to obtain a gamut of models with different trade-offs between complexity and performance by controlling the block diagonal structure size in the MLP.
+Vision Transformers have achieved impressive performance in many vision tasks.
+While the token mixer or attention block has been studied in great detail, much less research has been devoted to the channel mixer or feature mixing block (FFN or MLP), which accounts for a significant portion of the model parameters and computation.
+In this work, we show that the dense MLP connections can be replaced with a sparse block diagonal MLP structure that supports larger expansion ratios by splitting MLP features into groups.
+To improve the feature clusters formed by this structure we propose the use of a lightweight, parameter-free, channel covariance attention (CCA) mechanism as a parallel branch during training.
+This enables gradual feature mixing across channel groups during training whose contribution decays to zero as the training progresses to convergence.
+As a result, the CCA block can be discarded during inference, enabling enhanced performance at no additional computational cost.
+The resulting Scalable CHannEl MixEr (SCHEME) can be plugged into any ViT architecture to obtain a gamut of models with different trade-offs between complexity and performance by controlling the block diagonal MLP structure.
 This is shown by the introduction of a new family of SCHEMEformer models.
-Experiments on image classification, object detection, and semantic segmentation, with different ViT backbones, consistently demonstrate substantial accuracy gains over existing designs, especially under lower FLOPs regimes.
-For example, the SCHEMEformer establishes a new SOTA of 79.7% accuracy for ViTs using pure attention mixers on ImageNet-1K at 1.77G FLOPs.
+Experiments on image classification, object detection, and semantic segmentation, with 12 different ViT backbones, consistently demonstrate substantial accuracy/latency gains (upto 1.5% /20%) over existing designs, especially for lower complexity regimes.
+The SCHEMEformer family is shown to establish new Pareto frontiers for accuracy vs FLOPS, accuracy vs model size, and accuracy vs throughput, especially for fast transformers of small size.
 1 Introduction
-Figure 1: Left: Accuray-FLOPs tradeoff for various SCHEME models (denoted in blue) based on popular ViT architectures.
-Right: Accuracy-FLOPs trade-off of the proposed SCHEMEformer family derived from the Metaformer-PPAA-S12 model [57] with higher expansion ratios in the MLP blocks.
-The SCHEMEFormer family outperforms the baseline for the same FLOPs, with larger gains for smaller models.
-[…]
+Refer to caption Refer to caption Refer to caption
+(a) Accuracy-FLOPs (b) Accuracy-Parameters (c) Accuracy-Throughput
+Figure 1: Comparison of the proposed SCHEMEformer family, derived from the Metaformer-PPAA-S12 model [55] with higher expansion ratios in the MLP blocks, and many SOTA transformers from the literature.
+The SCHEMEFormer family establishes a new Pareto frontier (optimal trade-off) for a) accuracy vs.
+FLOPs, b) accuracy vs model size, and c) accuracy vs, throughput.
+SCHEMEformer models are particularly effective for the design of fast transformers (throughput between 75 and 150 images/s) with small model size.
+See Appendix Figure 9 for zoomed version.
+Vision Transformers (ViTs) [11, 28, 44, 27] are now ubiquitous in computer vision.
+They decompose an image into a set of patches which are fed as tokens to a transformer model [43] of two main components: a spatial attention module, which reweighs each token according to its similarity to the other tokens extracted from the image, enabling information fusion across large spatial distances, and a channel mixer that combines the feature channels extracted from all patches using a multi-layer perceptron (MLP or FFN).
+A bottleneck of this model is the quadratic complexity of the attention mechanism on the number of patches.
+Numerous ViT variants have been proposed to address the problem, using improved attention mechanisms or hybrid architectures that replace attention or combine it with convolutions.
+Much less research has been devoted to the channel mixer.
+Most models simply adopt the two-layer MLP block of [43], where channels are first expanded by a specified expansion ratio and then compressed to the original dimension.
+This is somewhat surprising since the mixer is critical for good transformer performance.
+For example, it is known that 1) pure attention, without MLPs or residual connections, collapses doubly exponentially to a rank one matrix [10], 2) training fails to converge without residual connections or MLP [55], and 3) replacing MLPs with more attention blocks (both spatial and channel attention) of equivalent computational complexity lowers the transformer accuracy [8].
+All these observations indicate that the channel mixer is an indispensable ViT component.
+In this work, we uantify how much the channel mixer module contributes to ViT performance and investigate how to improve the trade-off between complexity and accuracy of the ViT model.
+We show that enhanced design of the channel mixer can lead to significant improvements in transformer performance by introducing a novel Scalable CHannEl MixEr (SCHEME) that enables the design of models with larger expansion ratios.
+SCHEME is a generic channel mixer that can be plugged into existing ViT variants to obtain effective scaled-down or scaled-up model versions.
+We replace the channel mixer of a state of the art architecture (SOTA) for low-complexity transformers, the MetaFormer-PPAA-S12 [55], with SCHEME to obtain a new family of SCHEMEformer models with improved accuracy/complexity trade-off.
+This is illustrated in Figure 1, where the SCHEMEformer family is demonstrated to establish new Pareto frontiers for accuracy vs FLOPS, accuracy vs model size, and accuracy vs throughput, showing that SCHEME allows fine control over all these variables, while guaranteeing SOTA performance.
+These properties are shown to hold for image classification, object detection and semantic segmentation tasks, as well as for different architectures, such as T2T-ViT [57], CoAtNet [5], Swin Transformer [27], CSWin [9] and DaViT [8].
+To develop SCHEME, we start by studying the impact of the mixer channel expansion.
+Transformer performance is shown to increase with expansion ration until it saturates for an expansion ratio beyond 8.
+However, because mixer (MLP) complexity also increases with the dimension of the intermediate representation, naive channel scaling with expansion ratios larger than 4 causes an explosion of parameters and computation, leading to models of large complexity and prone to overfit.
+To achieve a better trade-off between dimensionality and computation, we leverage recent findings about the increased hardware efficiency of mixers with block diagonal structure [1, 6].
+These group the input and output feature vectors of a layer into disjoint subsets and perform matrix multiplications only within each group, as illustrated in Figure 2.
+We denote the resulting MLP block as Block Diagonal MLP (BD-MLP).
+Despite the lack of feature mixing across groups, we find that a transformer model equipped with the BD-MLP and a larger expansion ratio (to match both the parameter count and computation) achieves comparable or slightly higher accuracy than a baseline with dense MLP.
+This suggests that the lower accuracy of the block diagonal operations is offset by the gains of larger expansion ratios.
+Further analyzing the features learned by the different groups of the BD-MLP, we observe that they form feature clusters of similar ability to discriminate the target classes.
+To learn better feature clusters, we seek a mechanism capable of restoring inter-group feature communication during training without increasing parameters.
+For this, we propose a channel attention branch that reweighs the input features of the BD-MLP according to their covariance matrix, as illustrated in Figure 2.
+This attention mechanism is denoted as the channel covariance attention (CCA) block.
+The re-weighted features are then fused with the BD-MLP output by means of a weighted residual addition with learned weights $(\alpha,1-\alpha)$.
+As shown in Figure 2, SCHEME combines the sparse block diagonal structure of the BD-MLP, and the parameter-free CCA attention module, to implement a channel mixer extremely efficient in terms of parameters.
+Ablations of the evolution of the fusion weight $1-\alpha$ learned in the CCA branch (see Fig.
+2) over training show that it gradually decays to zero during training.
+This happens consistently across all layers of the model and across model architectures.
+Hence, while the CCA is important for the formation of good feature clusters during training, it can be removed at inference without any loss, as illustrated in Table 3 and Figure 2.
+As a result, the model accuracy improves over simply using the BD-MLP mixer, but inference complexity does not.
+This leads to an extremely efficient inference setup, both in terms of parameters and FLOPs.
+Overall, the paper makes the following contributions,
 Figure 2: Proposed SCHEME channel mixer.
@@
 Experiments show that CCA helps learn better feature clusters, but is not needed once these are formed.
-Vision Transformers (ViTs) [15, 33, 51, 32] have become ubiquitous in computer vision in recent years.
-Under this architecture, an image is decomposed into a set of patches which are considered as tokens and fed to a transformer model [49] of two main components: a spatial attention module, which reweighs each token according to its similarity to the other tokens extracted from the image, enabling information fusion across large spatial distances, and a channel mixer that combines the feature channels extracted from all patches using a multi-layer perceptron (MLP or FFN).
-A bottleneck of this model is the quadratic complexity of the attention mechanism on the number of patches.
-Numerous ViT variants have been proposed in recent years to improve complexity and/or accuracy.
-These can be grouped into 1) improved attention mechanisms, and 2) hybrid architectures that replace attention or combine it with convolutions.
-Much less research has been devoted to the channel mixer.
-Most models simply adopt the MLP block of [49], consisting of two layers, where channels are first expanded by a specified expansion ratio and then compressed to the original dimension.
-This is somewhat surprising since the mixer is critical for good transformer performance.
-For example, Dong et.
-al.
-[14] showed that pure attention, without MLPs or residual connections, collapses doubly exponentially to a rank one matrix.
-Yu et.
-al.
-[57] also showed that without residual connections or MLP, the model cannot converge.
-DaViT [12] showed that replacing MLPs with more attention blocks (both spatial and channel attention) of equivalent computational complexity lowers the transformer accuracy.
-These observations show that the channel mixer is an indispensable component of ViTs.
-In this work, we investigate the channel mixer module, both to quantify its contribution to ViT performance and to improve the complexity vs.
-accuracy trade-off of the transformer model.
-We start by studying the impact of the channel expansion performed in the mixer, by varying its expansion ratio.
-Transformer performance is shown to increase until it saturates for an expansion ratio beyond 8.
-However, because mixer (MLP) complexity is linear on the dimension of the intermediate representation, naive channel scaling with expansion ratios larger than 4 causes an explosion of parameters and computation, leading to models of large complexity and prone to overfit.
-To achieve a better trade-off between dimensionality and computation, we seek to leverage a hardware efficient mixer with block diagonal structure [1, 10].
-This consists of breaking the input and output feature vectors of a network layer into disjoint groups and performing matrix multiplications only within each group, as illustrated in Figure 2.
-We denote the resulting MLP block as Block Diagonal MLP (BD-MLP).
-Despite the lack of feature mixing across groups, we find that a transformer model equipped with the BD-MLP and a larger expansion ratio (to match both the parameter count and computation) achieves comparable or slightly higher accuracy than a baseline with the dense MLP.
-This suggests that the lower accuracy of the sparse block diagonal operations is offset by the gains of larger expansion ratios.
-Further analyzing the features learned by the different groups of the BD-MLP, we observe that they form feature clusters that are similar in terms of their discriminative ability to identify the target classes.
-To learn better feature clusters, we seek a mechanism capable of restoring inter-group feature communication only during training without increasing parameters and computational overhead.
-For this, we propose to use the channel correlations among features.
-This is done by introducing a channel attention branch that computes the channel covariance matrix of the input features of the BD-MLP and is used to re-weigh its output features, as illustrated in Figure 2.
-This attention mechanism is denoted as the channel covariance attention (CCA) block.
-The reweighted features are then fused with the BD-MLP output by means of a weighted residual addition with learned weights $(\alpha,1-\alpha)$.
-The overall design, combining the sparse block diagonal structure of the BD-MLP, and the parameter-free CCA attention module, is extremely efficient in terms of parameters.
-Experiments also show that the fusion weight learned in the CCA branch, denoted as $1-\alpha$ in Fig.
-2, gradually decays to zero during training.
-This happens consistently across all layers of the model and across model architectures.
-While the CCA is important for the formation of good feature clusters during training, it can be removed at inference without any loss, as illustrated in Figure 2.
-As a result, the model accuracy improves over simply using the BD-MLP mixer, but the computational complexity of model inference does not.
-This leads to an extremely efficient inference setup, both in terms of parameters and FLOPs.
-We empirically verify the hypothesis through ablations that study the evolution of attention branch weights over training and the class separability of the learned feature groups.
-Since the combination of BD-MLP and CCA enables the design of models with larger expansion ratios, we denote the resulting mixer as the Scalable CHannEl MixEr (SCHEME).
-SCHEME is a generic channel mixer that can be plugged into existing ViT variants to realize larger expansion ratios and obtain effective scaled-down or scaled-up model versions.
-We replace the channel mixer of the MetaFormer-PPAA-S12 [57] model with SCHEME to obtain a new family of SCHEMEformer models with improved accuracy/complexity trade-off.
-This is shown in the right of Figure 1.
-For the same parameter count and FLOPs, SCHEME obtains better accuracy than the baseline by using larger MLP expansion ratios.
-The gains are particularly larger for the smaller models.
-For example, the SCHEMEformer family achieves SOTA results of 79.74% accuracy at 1.77 G FLOPs, for ViTs using self-attention mixers.
-The left of Figure 1 shows that SCHEME models based on popular ViT architectures have a better Accuracy-FLOPs trade-off than the baseline models.
-Most of the SCHEMEformer models lie on the Pareto frontier of accuracy vs complexity, showing that SCHEME allows fine control over the two variables, while guaranteeing SOTA performance.
-These properties are shown to extend to object detection and semantic segmentation tasks, as well as to different architectures, such as T2T-ViT [60], CoAtNet [9], Swin Transformer [32], CSWin [13] and DaViT [12].
-In summary, our contributions are as follows,
-  • We study the channel mixer of MLPs in ViT and show that the dense feature mixing can be replaced by sparse feature mixing with higher internal feature dimensionality for improved accuracy.
-  • We propose the SCHEME module, combining 1) a BD-MLP to enable internal feature representations of larger dimensionality than previous MLP blocks, and 2) CCA to enable the learning of these representations without cost at inference.
+  • a study of the channel mixer of ViT MLPs, showing that dense feature mixing can be replaced by sparse feature mixing of higher internal feature dimensionality for improved accuracy, without increased complexity.
+  • the SCHEME module, which combines 1) a BD-MLP to enable internal feature representations of larger dimensionality than previous MLP blocks, and 2) CCA to enable the learning of these representations without cost at inference.
+  • various models that combine SCHEME with previous transformer architectures to achieve SOTA trade-offs between accuracy and model size, FLOPS, or throughput, such as the SCHEMEformer of Figure 1.
+This is shown particularly effective for the design of fast transformers with small model size, of interest for edge devices, robotics, and low-power applications.
   • SCHEME is shown to be generic and easily plugged into existing ViTs to obtain a gamut of models with improved performance.
 We use it to introduce a new family of SCHEMEformer models with better accuracy/complexity trade-off than previous ViTs.
-  • Experiments on image classification, object detection, and semantic segmentation consistently show gains in accuracy for fixed computation and model size.
+  • Experiments on image classification, object detection, and semantic segmentation, showing consistent gains in accuracy for fixed computation and size.
 2 Related Work
-Vision Transformers: Vision transformers advanced the state-of-the-art in several vision tasks since [33, 15] successfully applied the transformer-based self-attention NLP model of [49] to image generation and classification tasks.
-These models rely on a spatial attention mechanism, based on the matrix of dot-products between features extracted from image patches.
+Vision Transformers: Vision transformers advanced the SOTA in several vision tasks since [28, 11].
 This has quadratic complexity in the number of patches and is quite intensive.
-Most follow up work [19, 40, 59, 45, 41, 54, 43, 60, 62, 50] improved the spatial attention mechanism of ViT.
-DeiT [43] and subsequent works [46, 47] introduced a distillation token to distill information, typically from a CNN teacher, without large amounts of data or compute.
-PvT [51] proposed a progressive shrinking pyramid architecture with spatial-reduction attention that scales ViTs for dense prediction tasks beyond image classification.
-Swin transformers [32] introduced a hierarchical shifted window attention mechanism, which reduces the complexity to linear with respect to the number of windows.
-HaloNet [50] proposed two extensions for local ViTs [34] with blocked local attention and relaxed translational equivariance for scaling ViTs.
+Most follow up work [17, 35, 56, 40, 36, 48, 37, 57, 58, 42, 38, 39, 44, 27, 42, 29, 61, 3, 18, 15] improved the spatial attention mechanism of ViT.
 A more extensive review of ViTs is given in [26].
-Recently, several works have shown that spatial attention is not the critical ViT feature.
+Role of attention: Recently, several works have shown that spatial attention is not the critical ViT feature.
 Some works improved performance by relying on hybrid architectures, which augment or replace ViT layers with convolutions [47, 41].
-Efficient transformer designs, for edge devices, frequently sacrifice spatial attention to achieve better trade-offs between FLOPs and accuracy [5, 27].
 Other works have questioned the need for spatial attention altogether.
-Metaformer [57] argued that the fundamental trait of ViT is the mixing of information across patches, showing that competitive results can be obtained by simply replacing attention with pooling or identity operations.
-Similarly, [30] showed that spatial attention is not critical for vision transformers by proposing a spatial-gating MLP of comparable performance to ViT [15].
-DaViT [12] showed that it is helpful by simply reusing the spatial attention design for channel attention, building a cascade of alternating spatial and channel attention blocks.
-While this improves performance, it increases the complexity of the transformer block, resulting in a model with many parameters and potentially redundant channel mixer operations.
-Despite all this work on ViT architectures, little emphasis has been devoted to the other main component of the transformer: the channel mixer module (MLP) that follows attention.
+[25] showed that spatial attention is not critical for vision transformers by proposing a spatial-gating MLP of comparable performance to ViT.
+Similarly, Metaformer [55] argued that the fundamental trait of ViT is the mixing of information across patches, showing that competitive results can be obtained by simply replacing attention with pooling or identity operations.
+DaViT [8] enhances performance by cascading spatial and channel attention blocks but increases transformer complexity, adding parameters and redundant channel mixing operations.
+Role of MLP: Despite all this work on ViT architectures, little emphasis has been devoted to the channel mixer module (MLP) that follows attention.
 This is surprising because the mixer dominates both the parameter count and complexity (FLOPs) of the standard transformer block.
@@
 This improved performance but increases parameter and computation costs.
-Switch Transformer [17] replaces the FFN with sparse mixture of experts [39] that dynamically routes the input tokens.
-This design allows scaling models to large sizes using higher number of experts, however its not effective for ViTs.
-In this work, we instead propose an efficient and generic channel mixer module that improves both the parameter and computational efficiency of the transformer and allows for flexible scaling of ViTs.
+Switch Transformer [14] replaces the FFN with a sparse mixture of experts [34] that dynamically routes the input tokens.
+This design allows scaling models to large sizes using higher number of experts but is not effective for ViTs.
+XCiT [12] employs a cross-covariance attention (XCA) operator, which can be seen as a "transposed" version of self-attention that operates across feature channels.
+However, it still employs the standard feedforward network after the XCA operation.
+Some recent works [52, 13] focus on MLP modules either by increasing the non-linearity with an arbitrary GeLU function or including modules such as SwiGLU, LayerNorm and RoPE which increases the complexity and latency of the model.
+In this work, we augment the FFN or MLP with a modified Cross-Covariance Channel Attention (CCA) operation during training to facilitate feature mixing across different channel groups that is applicable to various ViTs including XCiT to improve the computation-accuracy trade-off.
 3 The SCHEME module
 Figure 2 depicts the proposed SCHEME module for feature mixing in ViTs.
-As shown in the figure, the standard channel mixer consists of two MLP layers, which expand the dimensionality of the input features and then reduce it to the original size.
+As shown on the left, the standard channel mixer consists of two MLP layers, which expand the dimensionality of the input features and then reduce it to the original size.
 Let $\mathbf{x}\in\mathbb{R}^{d\times N}$ be the matrix containing the $N$ $d$-dimensional input feature vectors extracted from $N$ image patches.
@@
 $\mathbf{y}=\mathbf{W}_{2}\mathbf{z}+\mathbf{b}_{2}{\bf 1}^{T}_{N}$ (2)
-where $\mathbf{W}_{1}\in\mathbb{R}^{Ed\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times Ed},\mathbf{b}_{1}\in\mathbb{R}^{d},\mathbf{b}_{2}\in\mathbb{R}^{Ed},$ ${\bf 1}_{N}$ is the $N$-dimensional vector containing ones as all entries, $\sigma(.)$ is the activation function (typically GELU), and $E$ is an expansion factor, typically 4.
+where $\mathbf{W}_{1}\in\mathbb{R}^{Ed\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times Ed},\mathbf{b}_{1}\in\mathbb{R}^{Ed},\mathbf{b}_{2}\in\mathbb{R}^{d},$ ${\bf 1}_{N}$ is the $N$-dimensional vector containing ones as all entries, $\sigma(.)$ is the activation function (typically GELU), and $E$ is an expansion factor, typically 4.
+Table 1: MetaFormer-S12 [55] ImageNet-1K validation accuracy vs MLP expansion ratio ($E$). {*}: results by author code.
+Ratio FLOPs Top-1 Acc
+Model ($E$) (G) (%)
+Metaformer 8 4.14 81.6 (+0.6)
+Metaformer 6 3.35 81.8 (+0.8)
+Metaformer [55] 4 2.55 81.0 (+0.0)
+Metaformer 2 1.77 78.9 (-2.1)
+Metaformer 1 1.33 76.0 (-5.0)
+[…]
+Table 2: SCHEME parameters (P) and FLOPs (F). 44-e8 (12-e8) downscales (upscale) the MetaFormer model.
+SCHEME
+44-e8 12-e8
+S12 P (M) 12 21
+S12 F (G) 1.8 3.3
+S24 P (M) 21 40
+S24 F (G) 3.3 6.5
+S36 P (M) 55 59
+S36 F (G) 8.0 9.6
+Table 3: Impact of removing CCA branch during inference.
+CCA Params FLOPs Top-1 Acc
+SCHEME Used  (M) (G) (%)
+44-e8-S12 11.83 2.16 79.74
+44-e8-S12 11.83 1.77 79.72
+12-e8-S24 40.0 7.3 82.80
+12-e8-S24 40.0 6.5 82.76
+CoatNet-44-e8 17.80 3.83 80.70
+CoatNet-44-e8 17.80 3.42 80.68
+Swin-12-e8-T 36.93 7.00 81.69
+Swin-12-e8-T 36.93 5.89 81.69
 Table 3 details the impact of the mixer on overall transformer performance by evaluating the role of the expansion factor $E$ on the performance of Metaformer-PPAA-S12 [55] (Pooling, Pooling, Attention, Attention) architecture on ImageNet-1K.
@@
 The table also shows that these gains are not trivial.
-The MetaFormer-S18, which has no expansion ($E=1$) but more transformer layers and complexity comparable to that of the MetaFormer-S12 ($E=4$), has a significantly lower accuracy of 78.3%.
+The S18 model, which has no expansion ($E=1$) but more transformer layers and complexity comparable to that of the S12 model with $E=4$, has an accuracy $2.7$ points lower than the latter.
 In summary, for a given computation budget, it is beneficial to trade off transformer depth for dimensionality expansion in the channel mixer.
@@
 3.1 Scalable Channel Mixer (SCHEME)
-To efficiently realize the gains of higher intermediate MLP feature dimensionality, we propose the SCHEME module, which consists of a BD-MLP and CCA block where the CCA branch is only used during training.
-The following sections describe these in detail.
-Block Diagonal MLP (BD-MLP): Block diagonal matrices have been previously proposed to efficiently approximate dense matrices [1, 10].
+Block Diagonal MLP (BD-MLP): Block diagonal matrices have been previously used to efficiently approximate dense matrices [1, 6].
 In CNNs, group channel operations are frequently used to design lightweight mobile models with improved accuracy-computation trade-off [22, 2, 31].
@@
 $\mathbf{z}_{k}=\sigma(\mathbf{W}_{1,k}\mathbf{x}_{k}+\mathbf{b}_{1,k}{\bf 1}^{T}_{N})$ (3)
-$\mathbf{y}_{k}=\mathbf{W}_{2.k}\mathbf{z}_{k}+\mathbf{b}_{2,k}{\bf 1}^{T}_{N}$ (4)
-where $\mathbf{W}_{1}\in\mathbb{R}^{Ed/g_{1}\times d/g_{1}},\mathbf{W}_{2}\in\mathbb{R}^{d/g_{2}\times Ed/g_{2}},\mathbf{b}_{1}\in\mathbb{R}^{d/g_{1}},\mathbf{b}_{2}\in\mathbb{R}^{Ed/g_{2}}$ and $\mathbf{z}$ is decomposed into a set $\{\mathbf{z}_{k}\}_{k=1}^{G}$ where $\mathbf{z}_{k}\in\mathbb{R}^{d/G\times N}$, with $G=g_{1}$ in (3) and $G=g_{2}$ in (4).
+$\mathbf{y}_{k}=\mathbf{W}_{2,k}\mathbf{z}_{k}+\mathbf{b}_{2,k}{\bf 1}^{T}_{N}$ (4)
+where $\mathbf{W}_{1,k}\in\mathbb{R}^{Ed/g_{1}\times N/g_{1}},\mathbf{W}_{2,k}\in\mathbb{R}^{d/g_{2}\times Ed/g_{2}},\mathbf{b}_{1}\in\mathbb{R}^{Ed/g_{1}},\mathbf{b}_{2}\in\mathbb{R}^{d/g_{2}}$ and $\mathbf{z}$ is decomposed into a set $\{\mathbf{z}_{k}\}_{k=1}^{G}$ where $\mathbf{z}_{k}\in\mathbb{R}^{Ed/G\times N}$, with $G=g_{1}$ in (3) and $G=g_{2}$ in (4).
 Since the complexity of (3) is $g_{1}^{2}$ times smaller than that of (1) and there are $g_{1}$ groups, the complexity of the first MLP is $1/g_{1}$ times that of standard MLP.
@@
 For example, when $g_{1}=g_{2}=g$ this allows growing the expansion factor by a factor of $g$ without computational increase.
-Expansion FLOPs Top-1 Acc
-Model ratio ($E$) (G) (%)
-[…]
-[…]
-Metaformer-PPAA-S12 [57] 4 2.55 81.0 (+0.0)
-[…]
-[…]
-[…]
-Table 1: Effect of the MLP block expansion ratio ($E$) on ImageNet-1K validation accuracy for the MetaFormer-S12 [57] architecture.
-*{*}* denotes results obtained with the code made publicly available by the authors.
 […]
 $CCA(\mathbf{x})=\mbox{softmax}\left(\frac{\mathbf{x}\mathbf{x}^{T}}{\tau}\right)\mathbf{x}$ (5)
-where the softmax operation is applied across the matrix rows and $\tau$ is a smoothing factor.
-The output of the channel mixer block is obtained by the weighted summation of the BD-MLP and CCA branches, according to
-$\begin{split}\textbf{y}_{out}=\alpha\textbf{y}+(1-\alpha)CCA(\textbf{x})\end{split}$ (6)
+where the softmax operation is applied across the matrix rows and $\tau$ is a smoothing factor empirically set to 1 in our experiments.
+The output of the channel mixer block is the weighted summation of the BD-MLP and CCA branches
+$\begin{split}\textbf{y}_{out}=\alpha\textbf{y}+(1-\alpha)CCA(\textbf{x}),\end{split}$ (6)
 where $\alpha$ is a mixing weight learned across all samples. Various other design choices are discussed in Section 4.2.
-CCA as a Regularizer: The design choice of using a parameter free attention branch and a static weight $\alpha$ allows the model to form better feature clusters during training and gradually decay the contribution from CCA branch once the feature clusters are formed.
-This can be seen in Figure 3(a), which plots the value of the learned mixing weight $1-\alpha$ as a function of training epochs on ImageNet-1K, for all transformer layers.
+CCA as a Regularizer: The introduction of a parameter free attention branch and a learnable weight $\alpha$ allows the model to form better feature clusters during training and gradually decay the contribution from CCA branch once the feature clusters are formed.
+This can be seen in Figure 4, which plots the value of the learned mixing weight $1-\alpha$ as a function of training epochs on ImageNet-1K, for all transformer layers.
 These plots are typical of the behavior we observed with all transformer backbones and architectures we considered.
 Clearly, $1-\alpha$ starts with high to intermediate values, indicating that information flows through both branches of the mixer, but decays to $1-\alpha\approx 0$ as training converges.
-Hence, there is no degradation if the CCA branch is removed during inference.
+Hence, as shown in Table 3, there is no degradation if the CCA branch is removed during inference.
 This eliminates a substantial amount of computation during inference, leading to the the training and inference setup of Figure 2, where CCA is not used at inference.
@@
 While this hypothesis is not trivial to test, since (5) varies from example to example, we confirmed that using CCA during training enhances class separability, which likely reduces overfitting for large expansion ratios.
-See the results section for more details.
+See section 4.2 for more details.
+Computational Complexity: The complexity of the BD-MLP block is controlled by the group numbers $g_{1}$, $g_{2}$ and the expansion factor $E$, with a total cost $\mathcal{O}(Ed^{2}/g_{1}+Ed^{2}/g_{2})$, where $d$ is the channel dimension.
+The computational cost of CCA is $\mathcal{O}(Nd^{2})$ where $N$ is the number of tokens.
+Since CCA is not used during inference, it only adds to the computations during training.
+SCHEME provides a systematic framework that balances transformer width and depth by adjusting block size and expansion hyperparameters.
 3.2 The SCHEMEformer family
-The proposed SCHEME module can be plugged into any ViT architecture, enabling efficient control of model complexity via the mixer hyperparameters $g_{1}$, $g_{2}$, and $E$.
+The proposed SCHEME module enables efficient control of model complexity via the mixer hyperparameters $g_{1}$, $g_{2}$, and $E$.
 Table 3 shows that naively scaling down the ViT model by simply reducing $E$ causes a significant accuracy loss.
@@
 This is demonstrated by the introduction of a new family of models, denoted as SCHEMEformer, obtained by replacing the channel mixer of the Metaformer-PPAA [55] architecture with the SCHEME module.
-Two such configurations are shown in Table 2 where the naming follows the convention {model-name}-{$g_{1}$$g_{2}$}-e{$E$}.
-These are shown to achieve SOTA accuracy for lower computational budgets in the experiments section, where we also discuss results with other backbones.
-S12 S24 S36
-P F P F P F
-Model (M) (G) (M) (G) (M) (G)
-SCHEMEformer-PPAA-44-e8 12 1.8 21 3.3 55 7.96
-SCHEMEformer-PPAA-12-e8 21 3.3 40 6.5 59 9.6
-Table 2: Scaling MetaFormer-PPAA using SCHEME.
-Each column of the table reports to one baseline model (S12 to S36) of increasing size.
-P and F denote Parameters and FLOPs respectively.
-SCHEMEformer-PPAA-44-e8 (SCHEMEformer-PPAA-12-e8) models use configurations $g_{1}$, $g_{2}$, and $E$ that downscale (upscale) the baseline.
-3.3 Computational Complexity
-The complexity of the BD-MLP block is controlled by the group numbers $g_{1}$, $g_{2}$ and the expansion factor $E$, with a total cost $\mathcal{O}(Ed^{2}/g_{1}+Ed^{2}/g_{2})$, where $d$ is the channel dimension.
-The computational cost of CCA is $\mathcal{O}(Nd^{2})$ where $N$ is the number of tokens.
-Since CCA is not used during inference, it only adds to the computations during training.
-The SCHEME framework provides a systematic way to control the trade-off of transformer width vs depth, by controlling the block size and expansion hyperparameters.
+See Table 3 for two such configurations.
 4 Experimental Results
-In this section, we discuss an evaluation of the SCHEME module on Image Classification (ImageNet-1K [11]), Object Detection (COCO-17 [28]), and Semantic Segmentation (ADE-20K [65]) benchmarks.
-#Params FLOPs Top-1 Acc
-Model  (M) (G) (%)
-Attention Mixers
-Poolformer-S7 [57] 8.60 1.1 73.0
-PVT-Tiny [51] 13 1.9 75.1
-XCiT-T12 [16] 7 1.2 77.1
-Poolformer-S12 [57] 12 1.8 77.2
-MetaFormer-PPAA-11-e2-S12 [57] 12 1.8 78.9
-XCiT-T24 [16] 12 2.3 79.4
-ViT-B/16 [15] 86 17.6 79.7
-SCHEME-Poolformer-44-e8-S12 7.20 1.0 73.0
-SCHEME-Poolformer-12-e8-S12 16.7 2.6 78.5
-SCHEME-CSWin-44-e8-Tiny 15.6 3.0 79.2
-SCHEME-Swin-44-e8-Tiny 19.7 3.1 79.6
-SCHEMEformer-PPAA-44-e8-S12 12 1.77 79.7
-PVT-Small [51] 25 3.8 79.8
-QuadTree-B-b1 [41] 14 2.3 80.0
-Poolformer-S24 [57] 21.0 3.4 80.3
-MetaFormer-PPAA-S12 [57] 17 2.6 81.0
-PVT-Medium [51] 44 6.7 81.2
-Swin-Tiny [32] 29 4.5 81.3
-T2T-ViT t-14 [60] 22 6.1 81.7
-DeiT-B [43] 86 17.5 81.8
-CaiT-XS-24 [45] 27 5.4 81.8
-ViL-Small [61] 25 5.1 82.0
-XCiT-S12 [16] 26 4.8 82.0
-SCHEME-Poolformer-12-e8-S24 30.8 4.9 80.5
-SCHEME-T2T-ViT-44-e8-14 17.4 6.1 81.7
-SCHEME-Swin-12-e8-Tiny 36.9 5.9 81.7
-SCHEMEformer-PPAA-44-e8-S24 21 3.31 81.9
-SCHEMEformer-PPAA-12-e8-S12 21 3.35 82.0
-SCHEME-T2T-ViT-12-e8-14 27.7 8.1 82.1
-SCHEME-DaViT-44-e8-Tiny 19.7 3.8 82.1
-Focal-Tiny [56] 29 4.9 82.2
-CrossViT-Base [2] 105 21.2 82.2
-T2T-ViT-24[60] 64 14.1 82.3
-CPVT-Base [7] 88 17.6 82.3
-DaViT-Tiny [12] 23.0 4.3 82.8
-CSWin-Tiny [13] 23.0 4.3 82.8
-SCHEMEformer-PPAA-12-e8-S24 40 6.47 82.8
-SCHEME-DaViT-12-e8-Tiny 37.0 6.6 83.0
-SCHEME-CSWin-12-e8-Tiny 29.1 5.6 83.2
-XCiT-L24 [16] 189 36.1 82.9
-TNT-B [20] 66 14.1 82.9
-Swin-Small [32] 50 8.7 83.0
-ViL-Base [61] 56 13.4 83.2
-CSWin-Small [13] 35.0 6.9 83.6
-PVTv2-B5 [52] 82.0 11.8 83.8
-Focal-Base [56] 89.8 16.0 83.8
-SCHEMEformer-PPAA-44-e8-S36 55 7.96 83.2
-SCHEMEformer-PPAA-12-e8-S36 58 9.58 84.0
-Convolution+Attention Mixers
-CeiT-S [59] 24 4.8 82.0
-CvT-13 [53] 20 4.5 81.6
-CoAtNet-0 [9] 25 4.2 81.6
-PatchConv-S60 [44] 25 4.0 82.1
-CvT-21 [53] 32 7.1 82.5
-ViTAEv2-S [63] 19 5.7 82.6
-$*{}^{*}$ 17.8 3.4 80.7
-$*{}^{*}$ 24 4.1 81.6
-SCHEME-CAformer-44-e8-S18 19 2.92 82.4
-SCHEME-CAformer-12-e8-S12 24 3.61 82.9
-Table 3: Image Classification on ImageNet-1K dataset.
-Comparison with SOTA ViT models grouped by the type of token mixers and accuracy.
-Proposed SCHEMEformer-PPAA-44-e8(12-e8) models are shown in light blue and the baseline architectures are shown in light cyan.
-*{*}* denotes our implementation as their code is not publicly available.
-Figure 3: Comparison of various SCHEMEformer (blue stars) and SOTA ViT architectures (orange circles) (calculated on NVIDIA-Titan-X GPU with batch size 1).
-Left: Accuracy-Throughput trade-off.
-Right: Accuracy-Parameters trade-off.
+Figure 3: Impact of CCA (SCHEMEformer-44-e8-S12).
+Left: Evolution of weight $1-\alpha$ across model layers.
+Right: Class separability of output features $\bf y$ (over 50 random classes of ImageNet-1K validation set) for model trained with and without CCA.
+See Appendix Figure 11 for zoomed version.
+Figure 4: SCHEME tradeoffs.
+Left: Accuracy vs FLOPs of various SCHEME models with different MLP configurations.
+Right: SCHEME mixer improves accuracy for fixed throughput or vice-versa for various popular ViT architectures.
+See Appendix Figure 11 for zoomed version.
 4.1 Comparisons to the state of the art
-Implementation Details: For image classification, we use a batch size of 1024, adjusted according to memory availability, and train all models on 8 GPUs with AdamW optimizer and cosine annealing learning rate schedule, of learning rate 1e-3.
-Warmup learning is applied in the first 5 epochs.
-Object detection is based on the setting of [57], where all backbones are trained with RetinaNet [29] and Mask-RCNN [22] detection heads, for a fair comparison with prior methods.
-A 1x learning schedule (12 epochs with AdamW optimizer and initial learning rate of 1e-4) is used and all models are implemented with the mmdetection [3] codebase, capping input sizes at 800, 1333 for the shorter and longer sides, respectively.
-For semantic segmentation, following prior methods [4, 24, 57], we train with a batch size of 16 for 80,000 iterations with AdamW optimizer and an initial learning rate of 2e-4.
-We use the mmsegmentation [8] codebase with an input size $512\times 512$ for training and resize the shorter side to 512 pixels for testing.
-For both detection and segmentation, imagenet pretrained models are used to initialize the backbones.
-#Par $AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$ RetinaNet 1× ResNet-18[21] 21.3 31.8 49.6 33.6 16.3 34.3 43.2 PoolFormer-S12[57] 21.7 36.2 56.2 38.2 20.8 39.1 48.0 ResNet-50[21] 37.7 36.3 55.3 38.6 19.3 40.0 48.8 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 PoolFormer-S24[57] 31.1 38.9 59.7 41.3 23.3 42.1 51.8 ResNet-101[21] 56.7 38.5 57.8 41.2 21.4 42.6 51.1 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 Mask R-CNN 1× #Par $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ Mask R-CNN 1× ResNet-18[21] 31.2 34.0 54.0 36.7 31.2 51.0 32.7 PoolFormer-S12[57] 31.6 37.3 59.0 40.1 34.6 55.8 36.9 ResNet-50[21] 44.2 38.0 58.6 41.4 34.4 55.1 36.7 SCHEMEformer-44-e8-S12 31 39.8 61.9 42.9 24.1 53.0 42.3 PoolFormer-S24[57] 41.0 40.1 62.2 43.4 37.0 59.1 39.6 ResNet-101[21] 63.2 40.4 61.1 44.2 36.4 57.7 38.8 SCHEMEformer-44-e8-S24 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 4: COCO-17 Object Detection and Instance Segmentation.
-All backbones are pretrained on ImageNet-1K and use 1x learning schedule.
-$AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
-Backbone #Par FLOPs mIoU (%) ResNet-18[21] 16 32.2 32.9 PVT-Tiny[51] 17 33.2 35.7 ResNet-50[21] 29 45.6 36.7 PoolFormer-S12[57] 16 30.9 37.2 ResNet-101[21] 48 65.1 38.8 ResNeXt-101-32x4d[55] 47 64.7 39.7 PVT-Small[51] 28 44.5 39.8 XCiT-T12/8[16] 8.4 - 39.9 ResNeXt-101-64x4d[55] 86 103.9 40.2 PoolFormer-S24[57] 23 39.3 40.3 SCHEMEformer-44-e8-S12 15.5 34.3 40.9 PVT-Medium[51] 48 61.0 41.6 PoolFormer-S36[57] 35 47.5 42.0 PVT-Large[51] 65 79.6 42.1 PoolFormer-M36[57] 60 67.6 42.4 SCHEMEformer-44-e8-S24 24.8 45.7 42.5 Table 5: Semantic Segmentation results on ADE20K dataset using Semantic FPN [24].
-All backbones are pretrained on ImageNet-1K.
-FLOPs are calculated at $512\times 512$ resolution.
-Model BDM CCA Acc (%) Baseline 78.9 44-e8-S12 79.1 (+0.2) 44-e8-S12 79.7 (+0.6) Table 6: Ablation study of the contribution of BD-MLP and CCA branches to the overall performance for SCHEMEformer-PPAA-44-e8-S12 model.
-Baseline refers to Metaformer-PPAA-11-e2-S12.
-Module #Par FLOPs Top-1  (M) (G) Acc (%) Shuffle 11.83 1.77 79.1 SE 11.98 1.77 79.3 Conv 13.31 1.97 79.6 DyCCA 11.90 1.83 79.6 CCA 11.83 1.77 79.7 Table 7: Ablation study of alternative designs for feature attention branch in SCHEME.
 Image Classification: Image classification is evaluated on Imagenet-1K, without using extra data.
 We report the results with single crop top-1 accuracy at $224\times 224$ input resolution.
-Results are presented for various backbones and hyperparameter configurations, to show the generalizability of the SCHEME module.
-We evaluate the SCHEMEformer family of models based on the Metaformer-PPAA-S12 [57] obtained by replacing the MLP of the original model with the SCHEME module.
-Additionally, we replace the pooling of the first two stages of the Metaformer-PPAA-S12 with convolution  [58] to evaluate SCHEME-CAformer models.
-Table 3 compares SCHEME models to various SOTA models from the literature grouped by the type of token mixers and accuracy.
-The SCHEME models are highlighted in blue and the baseline models on which they are based in cyan.
-While the table is ordered by increasing accuracy, meaningful comparisons must account for both accuracy and parameter size or FLOPs, since the models have very different sizes.
-The SCHEMEformer family achieves the best trade-off between accuracy and computation or parameter size by using a larger internal feature dimension for the MLP where the gains are larger for smaller models, whose performance is not near saturation.
-For example, as highlighted in red font, for the same compute and parameter cost, the SCHEMEformer-PPAA-44-e8-S12 improves on the performance of the baseline MetaFormer-PPAA-11-e2-S12 by +0.8% and achieves a SOTA result of 79.7% top-1 accuracy at 1.77G FLOPs.
-The table also shows the accuracy of the SCHEME module applied to other popular ViT architectures, such as T2T-ViT, Swin, CSWin, Da-ViT in the Attention Mixers category, and CoatNet, CAformer in the Convolution+Attention Mixers category realizing models of intermediate complexity with comparable or better accuracy for the same parameters and FLOPs.
-For example, SCHEME-CoAtNet-12-e8-0 achieves 81.6% accuracy with 1M fewer parameters and 0.1G fewer FLOPs than the baseline CoAtNet-0 (highlighted in brown).
-SCHEME-CSWin-12-e8-Tiny (blue) achieves 83.2% accuracy with just 29 M parameters and 5.6 G FLOPs sandwiched between CSWin-Tiny (82.8%) and CSWin-Small (83.6%).
-Figure 1 provides a visual illustration of the results in the table.
-The plot on the left compares the SCHEME family (blue stars) to all other models (orange dots) in terms of the trade-off between FLOPS and accuracy.
-It is clear that most SCHEME models lie on the Pareto frontier for these two objectives, significantly expanding the set of models on this frontier.
+We evaluate the SCHEMEformer family of models based on the Metaformer-PPAA-S12 [55] obtained by replacing the MLP of the latter with the SCHEME module.
+Refer to Appendix A.2 for implementation details.
+We start by evaluating how this improves the trade-off between model accuracy and complexity.
+As discussed in Section 3.1, when $g_{1}=g_{2}$, a SCHEME transformer of expansion factor $gE$ has identical complexity to a standard transformer of expansion factor $E$.
+Hence, for fixed FLOPS, SCHEME allows an increase of the expansion factor by $g$.
+Figure 4 a) compares the performance of the Metaformer-PPAA-S12 with expansion ratios $E\in\{1,2,4,8\}$ to comparable variants of the SCHEMEformer-PPAA-S12, with SCHEME mixers of either $g=2$ (green curve) or $g=4$ (blue curve) groups.
+The SCHEMEformer models have a better trade-off between accuracy and FLOPS, achieving higher accuracies for all complexity levels.
+Among these, the one with more feature groups ($g=4$) has the best performance.
+While SCHEMEformer gains are observed for all FLOP levels, they are larger for lower complexity models.
+This makes SCHEME particularly attractive for the design of low complexity transformers, e.g.
+for edge devices.
+We next compare the SCHEMEformer family against the SOTA transformers in the literature.
+This is not an easy comparison, since models vary in size, FLOPS, and throughput.
+Because it is difficult to make any of these variables exactly the same for two different architectures, the comparison is only possible in terms of how the different architectures trade-off accuracy for any of the other factors.
+For a given pair of variables, e.g.
+FLOPS vs accuracy, the model is said to be on the Pareto frontier of the two variables if it achieves the best trade-off between the two.
+Table 4.1 presents a comparison of the SCHEMEformer family against various SOTA transformers in the literature.
+Each section of the table compares a SCHEMEformer model to a group of SOTA transformers of equivalent size or complexity.
+Note that, in each section, the remaining models have both lower throughput and accuracy than the SCHEMEformer model.
+In many cases they also have more parameters and FLOPs.
+Figure 1 provides a broader visualization of how SCHEME models establish Pareto frontiers for accuracy vs FLOPS, accuracy vs throughput, and accuracy vs model size (parameters).
+Figure 1 a) illustrates the trade-off between accuracy and FLOPS of many SOTA transformers.
+The dashed line connects the SCHEMEformer model results, summarizing the accuracy-FLOPs trade-off of the family.
+It can be seen that the SCHEME models lie on the Pareto frontier for these two objectives.
 This illustrates the fine control that SCHEME allows over the accuracy/complexity trade-off of transformer models.
-The plot on the right shows how the SCHEMEFormer models improve on the accuracy/complexity trade-off of the MetaFormer model on which they are based.
-This shows that the SCHEME module generalizes well to different backbones and offers higher flexibility in scaling.
-Throughput and Memory: Beyond higher accuracy, SCHEMEformer models also have low latency and parameter counts.
-The left of Figure 3 shows a favorable throughput vs accuracy trade-off for various SCHEME models compared to SOTA ViT architectures.
-Throughput is computed on NVIDIA-Titan-X GPU with a batch size of 1 averaged over 1000 runs.
-Despite having larger internal dimensions, the SCHEMEformer-PPAA-12-e8-S12 has a throughput (130 images per second) comparable to that of the baseline MetaFormer-PPAA-11-e2-S12.
-This is a consequence of the block diagonal structure, which enables higher FLOPs utilization [18].
-The right of Figure 3 shows a favourable accuracy vs parameters tradeoff comparison for SCHEME models compared to SOTA ViT architectures.
-SCHEME based scaling achieves higher accuracy with lower parameter counts.
-These results show that SCHEME models lie on the Pareto frontiers of accuracy vs throughput and vs memory consumption.
+Figure 1 b) presents a similar comparison for model sizes.
+Like for FLOPS, the SCHEME models lie on the Pareto frontier for accuracy vs model size.
+In fact the two plots are quite similar, showing that in general there is a good correlation between model size and FLOPS.
+Table 4: Image Classification on ImageNet-1K. Comparison of SCHEME models using expansion ratio 8 with SOTA ViTs grouped by accuracy. SCHEME family has higher throughput and accuracy than SOTA models.
+Model #Params FLOPs Thru Top-1 Acc
+ (M) (G) (im/s) (%)
+gMLP-Ti [25] 6 1.4 - 72.3
+ViT-L/16 [11] 307 63.6 37 76.1
+Meta-11-e2-S12 [55] 12 1.8 133 78.9
+MogaNet-T [23] 5 1.10 44 79.0
+XCiT-T24 [12] 12 2.3 - 79.4
+ViT-B/16 [11] 86 17.6 112 79.7
+SCHEME-44-S12 12 1.77 133 79.7
+$S^{2}$-MLP-deep [54] 51 10.5 - 80.7
+Meta-S12 [55] 17 2.6 87 81.0
+Swin-Tiny [27] 29 4.5 100 81.3
+T2T-ViT t-14 [57] 22 6.1 70 81.7
+DeiT-B [37] 86 17.5 114 81.8
+ViL-Small [59] 25 5.1 - 82.0
+SCHEME-12-S12 21 3.35 130 82.0
+Focal-Tiny [53] 29 4.9 29 82.2
+CPVT-Base [4] 88 17.6 - 82.3
+DaViT-Tiny [8] 23.0 4.3 61 82.8
+CSWin-Tiny [9] 23.0 4.3 20 82.8
+SCHEME-12-S24 40 6.47 69 82.8
+XCiT-L24 [12] 189 36.1 - 82.9
+Swin-Small [27] 50 8.7 31 83.0
+ViL-Base [59] 56 13.4 - 83.2
+CSWin-Small [9] 35 6.9 11 83.6
+SCHEME-12-S36 58 9.58 38 84.0
+Table 5: Semantic Segmentation results on ADE20K. FLOPs calculated at $512$ resolution for Semantic FPN and $1024$ resolution for UperNet.
+Backbone #P F mIoU
+(M) (G) (%)
+Semantic FPN
+ResNet-18[19] 16 32.2 32.9
+PVT-Tiny[44] 17 33.2 35.7
+ResNet-50[19] 29 45.6 36.7
+PoolFormer-S12[55] 16 30.9 37.2
+ResNet-101[19] 48 65.1 38.8
+ResNeXt-101-32x4d[51] 47 64.7 39.7
+PVT-Small[44] 28 44.5 39.8
+XCiT-T12/8[12] 8.4 - 39.9
+PoolFormer-S24[55] 23 39.3 40.3
+SCHEMEformer-44-S12 15.5 34.3 40.9
+PVT-Medium[44] 48 61.0 41.6
+PoolFormer-S36[55] 35 47.5 42.0
+PVT-Large[44] 65 79.6 42.1
+PoolFormer-M36[55] 60 67.6 42.4
+SCHEMEformer-44-S24 24.8 45.7 42.5
+UperNet
+Swin-Tiny [27] 60 945 44.5
+PVT-Large [45] 65 318 44.8
+Focal-Tiny [53] 62 998 45.8
+XCiT-S12/16 [12] 52 45.9
+DaViT-Tiny [8] 60 940 46.3
+SCHEME-DaViT-12-Tiny 68 969 47.1
+This is not the case for throughput, which for transformers is known to not necessarily correlate with FLOPS, due to GPU parallelism.
+For example, Table 4.1 shows that, CSWin models have lower throughput despite having lower FLOPs while ViT has higher throughput despite having higher FLOPs.
+SCHEMEformer controls this trade-off by controlling the expansion ratio and block diagonal structure, which enables higher FLOPs utilization for a given throughput [16].
+We compute throughput on a NVIDIA-Titan-X GPU with a batch size of 1 with input size 224x224 averaged over 1000 runs.
+While comparisons could be made for larger batch sizes, we consider the setting for live/streaming applications, where speed is most critical.
+In these applications, the concern is usually inference throughput, which requires batch size of 1.
+Figure 1 c) illustrates the trade-off between accuracy and throughput of various models, including very fast ResNet models of low accuracy.
+It can bee seen that the SCHEMEformer again achieves the best trade-off between these two variables, thus lying on the Pareto frontier for accuracy vs throughput.
+Its performance is particularly dominant in the range of throughputs between 75 and 150 images/sec, where it significantly outperforms the other methods.
+These results demonstrate how the SCHEME module endows transformer designers with the ability to produce multiple models at different points of the Pareto frontiers of accuracy vs model size, FLOPS, or throughput.
+SCHEME with other ViTs: Figure 4 shows the Accuracy-Throughput curves for various SCHEME models obtained by replacing the MLP blocks of popular ViT architectures.
+Table A in Appendix shows the detailed accuracy-latency comparisons of SCHEME with recent ViTs.
+The table shows the results for 12 different ViT backbones ranging from small (7.5 M) to large (143 M) models.
+It shows that SCHEME outperforms the baseline ViT models across on all 12 models either in accuracy or throughput.
+This demonstrates that SCHEME improves the accuracy for a fixed throughput or the throughput for a fixed accuracy, or both.
+The gains can be substantial, e.g.
+about 1.5% accuracy gain (constant throughput) for the fastest transformer or speed gains of up to 20% at constant accuracy.
+The table also shows that SCHEME is effective when scaling the model size as seen with SCHEME-ViT Large model.
+This shows that SCHEME benefits various ViT backbones and not just the MetaFormer.
+Semantic Segmentation: Table 4.1 compares the semantic segmentation performance of two SCHEMEformer models using the semantic FPN framework [21] to various SOTA models of similar complexity, on ADE20K.
+Since we do not not have access to the throughput of most models, we report only parameter sizes and FLOPS.
+In each section of the table, the remaining models have comparable or larger FLOPs and model size but lower accuracy than the corresponding SCHEMEformer.
+For example, SCHEMEformer-44-e8-S12 achieves 40.9% mIoU, which is 5.2/3.7 points higher than the PvT-Tiny/PoolFormer-S12, which both have comparable size and FLOPs.
+Similarly, the S24 model outperforms PoolFormer-M36 using only 41% of its parameters.
+To demonstrate the applicability of SCHEME to larger models, we also present a comparison of the SCHEME version of the DaViT-Tiny using the UperNet framework [50].
+While the DaViT-Tiny is already the best model in the table, the use of the SCHEME mixer improves its performance by an additional 0.8 points.
 Object Detection: Table 5 compares SCHEMEformer models to models of similar complexity on the COCO-17 object detection and instance segmentation benchmark, for both Retinanet and Mask-RCNN detection heads.
-The table is ordered by increasing mAP ($AP$, $AP^{b}$) values.
-Again, the SCHEMEformer models outperform all other models of the same or smaller size.
-For example, SCHEMEformer-44-e8-S12 with RetinaNet head achieves 38.3% overall mAP with just 21M params, outperforming models of comparable size by more than 2% and achieving comparable performance to models with $2.5$ times their size.
-Similar gains are obtained for the Mask-RCNN based SCHEMEformer-44-e8-S12.
-These results show that the SCHEME module also advances object detection.
-Semantic Segmentation: Table 5 compares SCHEMEformer-44-e8 S12 and S24 models using the semantic FPN framework [24] to other models of similar complexity on ADE20K semantic segmentation benchmark.
-As before, the SCHEMEformer models outperform all other models of the same or smaller size.
-For example, SCHEMEformer-44-e8-S12 achieves 40.9% mIoU, which is 5.2% higher than the model of most comparable size, PvT-Tiny.
-Similarly, the S24 model outperforms PoolFormer-M36 using only 41% of its parameters.
-In general, the gains obtained for semantic segmentation are higher than those of object detection, suggesting that the feature clusters formed by the block diagonal structure have larger benefits for dense prediction tasks.
+Again, the SCHEMEformer models outperform most other models of the same or smaller size.
+The only exception is the PoolFormer-S24, which slightly outperforms ($0.1$ points) the comparable SCHEME-former-44-S24, for the RetinaNet head.
+However, with the stronger Mask R-CNN head, the SCHEMEformer-44-S24 beats the PoolFormer-S24 by $0.8$ points.
+For the top performing models in the bottom third of each section of Table 5, SCHEME-DaViT-Tiny improves over DaViT-Tiny by an additional 0.7% and 0.9%, for RetinaNet and MaskRCNN heads respectively, while maintaining a comparable throughput.
+Table 6: COCO-17 Object Detection and Instance Segmentation.
+All backbones pretrained on ImageNet-1K (1x learning schedule).
+($AP^{b}$, $AP^{m}$): (bounding box AP, mask AP).
+$T$: throughput (images/sec).
+RetinaNet 1× #Par T $AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$
+PoolFormer-S12[55] 21.7 13.1 36.2 56.2 38.2 20.8 39.1 48.0
+ResNet-50[19] 37.7 15.7 36.3 55.3 38.6 19.3 40.0 48.8
+SCHEME-44-e8-S12 21 6.1 38.3 58.0 40.4 21.0 41.4 52.3
+PoolFormer-S24[55] 31.1 8.9 38.9 59.7 41.3 23.3 42.1 51.8
+ResNet-101[19] 56.7 12.1 38.5 57.8 41.2 21.4 42.6 51.1
+SCHEME-44-e8-S24 31 3.5 38.8 58.7 41.2 22.5 41.5 53.5
+DAT-T [49] 38 - 42.8 64.4 45.2 28.0 45.8 57.8
+CrossFormer-S [46] 41 - 44.4 55.3 38.6 19.3 40.0 48.8
+DaViT-Tiny [8] 39 8.2 44.0 65.6 47.3 29.6 47.9 57.3
+SCHEME-DaViT 47 7.8 44.7 66.2 48.3 30.0 48.8 57.2
+Mask R-CNN 1× #Par T $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+PoolFormer-S12[55] 31.6 9.9 37.3 59.0 40.1 34.6 55.8 36.9
+ResNet-50[19] 44.2 15.4 38.0 58.6 41.4 34.4 55.1 36.7
+SCHEME-44-e8-S12 31 6.0 39.8 61.9 42.9 37.1 59.2 39.4
+PoolFormer-S24[55] 41.0 7.9 40.1 62.2 43.4 37.0 59.1 39.6
+ResNet-101[19] 63.2 12.1 40.4 61.1 44.2 36.4 57.7 38.8
+SCHEME-44-e8-S24 41 3.4 40.9 62.5 44.6 37.8 59.7 40.4
+DAT-T [49] 48 - 44.4 67.6 48.5 40.4 64.2 43.1
+CrossFormer-S [46] 50 - 45.4 68.0 49.7 41.4 64.8 44.6
+DaViT-Tiny [8] 48 7.8 45.0 68.1 49.4 41.1 64.9 44.2
+SCHEME-DaViT 57 7.4 45.9 68.3 50.2 41.5 65.4 44.3
+Table 7: Contribution of BD-MLP and CCA branches in SCHEME.
+Model BDM CCA Acc (%)
+Baseline 78.9
+44-e8-S12 79.1
+44-e8-S12 79.7
+Table 8: Alternative feature attention designs in SCHEME.
+Module #Par FLOPs T Top-1
+ (M) (G) (img/s) Acc (%)
+Shuffle 11.83 1.77 108 79.1
+SE 11.98 1.77 86 79.3
+Conv 13.31 1.97 95 79.6
+DyCCA 11.90 1.83 75 79.6
+CCA 11.83 1.77 133 79.7
+Figure 5: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module.
+CCA G1 G2 G3 G4 Ensemble
+51.1 54.3 55.4 54.7 73.2
+47.6 50.0 49.2 55.1 73.8
 4.2 Ablation Studies
-In this section, we discuss ablations of the contribution of the BD-MLP and CCA branches, alternative designs for the channel mixer, and the regularizing effect of CCA.
-Contribution of BD-MLP and CCA branch: Table 6 shows an ablation of the contribution of the BD MLP and CCA branches to transformer performance.
-Starting from the Metaformer-PPAA-11-e2-S12, with expansion ratio $E=2$ and dense MLP, we replace the channel mixer with that of the SCHEME to obtain SCHEMEformer-PPAA-44-e8-S12, which has $g_{1}=g_{2}=4$ and $E=8$, maintaining the number of parameters and FLOPs fixed.
+Contribution of BD-MLP and CCA branch: Table 5 shows an ablation of the contribution of the BD-MLP and CCA branches.
+Starting from the Metaformer-PPAA-11-e2-S12, with expansion ratio $E=2$ and dense MLP, we replace the channel mixer by SCHEME to obtain the SCHEMEformer-PPAA-44-e8-S12 ($g_{1}$=$g_{2}$=4 and $E$=8), which maintains the number of parameters and FLOPs constant.
 The SCHEME model with only BD-MLP improves on the baseline by 0.2%.
@@
 Since CCA is not used at inference, its gains are free in terms of additional parameters/FLOPs.
-Effect of large expansion ratios: The right of Figure 1 shows the effect of simultaneously increasing the expansion ratio $E$ and adjusting groups to realize different models of similar size and complexity.
+Regularizing effect of CCA: Fig.
+4 shows the evolution of the weight $1-\alpha$ of the CCA branch in (6) during training.
+While initially large, it gradually decays to zero as training progresses.
+This holds for all network layers.
+Hence, CCA can be discarded at inference.
+Fig.
+1 in supplementary plots the weights $1-\alpha$ upon training convergence, for the family of SCHEMEformer-PPAA-44-e8 models, confirming that the weights are indeed very close to zero across all layers.
+Effect of large expansion ratios: The left of Figure 4 shows the effect of simultaneously increasing the expansion ratio $E$ and adjusting groups to realize different models of similar size and complexity.
 All models are based on the Metaformer-PPAA-S12.
@@
 The gains increase with larger expansions and saturate at larger FLOPs.
-This shows that higher internal feature dimensions are crucial for obtaining better accuracy with smaller ViT models.
+This shows that higher internal feature dimensions are important for obtaining better accuracy with smaller ViTs.
 Alternative designs of Channel Mixer: We investigate whether alternative choices to the CCA branch could accomplish this goal more effectively.
-Table 7 compares the performance of models that replace CCA with other feature mixing operations: the channel shuffling operation of ShuffleNet [64], a squeeze and excitation [23] network (SENet), a single layer of convolution, and a dynamic version of CCA (DyCCA), where the weight $\alpha$ of (6) is predicted dynamically, using GCT attention [35].
-The table shows that CCA obtains the best result.
-It is worth noting that while CCA has a lot more computation than some of these alternatives, it is not needed during inference.
-We found this not to be true for the alternatives, which produce much more balanced weights $\alpha$ after training convergence, and cannot be discarded at inference without a performance drop.
-We conjecture that, because all these alternatives have learnable parameters, the network learns to leverage them to extract complementary features, which must be fused with those produced by the BD-MLP branch during inference.
-Thus, while the alternative designs are very small modules, they incur an additional computational cost.
-CCA does not, achieving the best performance with no additional cost during inference.
-[…]
-Regularizing effect of CCA: Fig.
-3(a) shows the evolution of the weight $1-\alpha$ of the CCA branch in (6) during training.
-While initially large, it gradually decays to zero as training progresses.
-The figure shows that this holds for all network layers.
-Hence, CCA can be discarded at inference.
-Fig.
-5 in appendix plots the weights $1-\alpha$ upon training convergence, for the family of SCHEMEformer-PPAA-44-e8 models, confirming that the weights are indeed very close to zero across all layers.
-Table 8 shows the impact of removing CCA at inference in top-1 accuracy.
-The performance decrease is negligible ($\approx 0.02$) across backbones.
-[…]
-(a) Evolution of weight $1-\alpha$
-(b) Class separability
-Figure 4: Regularizing effect of CCA.
-Left: Evolution of weight $1-\alpha$ across all layers for SCHEMEformer-PPAA-44-e8-S12 model.
-Right: Comparison of class separability of the features $\bf y$ of the SCHEME module, for 50 random classes on ImageNet-1K validation set, using SCHEMEformer-PPAA-44-e8-S12 model trained with and without CCA.
-We conjectured above that CCA helps training because it facilitates feature clustering into naturally independent groups that do not require inter-group communication.
-We tested this hypothesis by studying the intermediate feature vectors ${\bf v}_{l}$ (obtained after max-pooling the features $\bf y$ of (4)) extracted from four randomly selected layers $l$ of two ImageNet pretrained models, trained with and without CCA respectively.
+Table 5 compares models that replace CCA with other feature mixing operations: the channel shuffling operation of ShuffleNet [60], a squeeze and excitation [20] network (SENet), a single layer of convolution, and a dynamic version of CCA (DyCCA), where the weight $\alpha$ of (6) is predicted dynamically, using GCT attention [30].
+CCA obtains the best result.
+While CCA is computationally heavier than some of these alternatives, it is not needed at inference, as shown in Table 3.
+This is not true for the alternatives, which produce much more balanced weights $\alpha$ after training convergence, and cannot be discarded at inference without performance drop.
+We conjecture that their learnable parameters extract complementary features which must be used at inference.
+Feature Clustering: We conjectured above that CCA helps training because it facilitates feature clustering into naturally independent groups that do not require inter-group communication.
+We tested this hypothesis by studying the intermediate feature vectors ${\bf v}_{l}$ (obtained after max-pooling the features $\bf y$ of (4)) extracted from four randomly selected layers $l$ of two ImageNet pretrained models, trained with and without CCA.
 We split the features into the 4 groups used in the model ${\bf v}_{l,g},g\in\{1,\ldots 4\}$ and concatenated the features of all layers in the same group.
@@
 Without CCA, i.e.
-no group communication during training, the network produces feature groups that tend to be individually more predictive of the image class, but less predictive when combined.
-This suggests that there is some amount of redundancy between the features extracted by the different groups.
-By introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
+no group communication during training, the network produces feature groups individually more predictive of the image class, but less predictive when combined.
+This suggests that there is redundancy between the features of the different groups.
+By introducing inter-group communication, CCA enables the groups to learn more diverse features, that complement each other.
 Fig.
-3(b) shows the class separability of the intermediate features $\bf y$ of (4) of a randomly chosen layer of the SCHEMEformer-PPAA-44-e8-S12 model.
+4 shows the class separability of the intermediate features $\bf y$ of (4) of a randomly chosen layer of the SCHEMEformer-PPAA-44-e8-S12.
 Class separability was measured as in [32], with a final value obtained by averaging the class separability across all classes.
@@
 Conversely, we tested if CCA is helpful when channel shuffling is inserted in between the two mixer MLP layers, which destroys the group structure.
-This variant of the SCHEMEformer-PPAA-44-e8-S12 model achieved an accuracy of 79.1% for both training with and without CCA.
+This variant of the SCHEMEformer-PPAA-44-e8-S12 model achieved an accuracy of 79.1% for both training with and without CCA (Table 5, row 1).
 This shows that CCA is not helpful when feature groups are mixed.
-Similarly, CCA did not provide any gains when applied to the standard MLP branch with full feature mixing (one group).
-All these results suggest that CCA indeed helps to form the independent feature clusters needed to achieve the computational efficiency of channel groups without performance degradation.
-5 Conclusion
-In this work, we proposed the SCHEME module with sparse feature mixing for improving the performance of ViTs.
-This uses a weighted fusion of a BD-MLP branch, which abstracts existing MLPs with block diagonal structure, and a parameter-free CCA branch that helps to cluster features into groups during training.
+Similarly, CCA did not provide any gains when applied to the standard MLP branch with full feature mixing.
+These results suggest that CCA indeed helps to form the independent feature clusters needed to achieve the computational efficiency of channel groups without performance degradation.
+Table 5 shows that, despite the higher accuracy of the individual group features of the model trained without CCA, the model trained with CCA has 0.6% higher ensemble accuracy.
+5 Limitation and Conclusion
+Although SCHEME module improves the accuracy-throughput curves of popular ViTs, it incurs a slight overhead in memory (see section A.3.2) during training due to the channel attention operation which limits the batch size on smaller GPUs.
+However, Figure 4 shows that the contribution of the CCA branch becomes negligible after 150 epochs.
+Therefore, the CCA branch can be removed beyond this point, significantly reducing the training overhead.
+In this work, we proposed the SCHEME module for improving the performance of ViTs.
+SCHEME leverages a block diagonal feature mixing structure to enable MLPs with larger expansion ratios, a property that is shown to improve transformer performance, without increase of model parameters or computation.
+During training, it uses a weighted fusion of a BD-MLP branch, and a parameter-free CCA branch that helps to cluster features into groups.
 The CCA branch was shown to improve training but not be needed at inference.
 Experiments showed that it indeed improves the class separability of the internal feature representation of the BD-MLP branch, helping create feature clusters that are informative of the image class.
-The standard MLP in transformer block is replaced with the SCHEME module to obtain a new family of SCHEMEformer models that were shown to improve performance for classification, detection, and segmentation, for fixed parameters and FLOPs with favorable latency.
-SCHEME is shown to be effective for various ViT architectures and was also shown to provide a flexible way to scale models, always outperforming models with smaller MLP expansion ratios having the same complexity.
-Acknowledgment
-We acknowledge and thank the use of the Nautilus platform for some of the experiments discussed above.
-Appendix
-Appendix A Ablation Studies
-CCA #Params FLOPs Top-1 Acc
-Model Used  (M) (G) (%)
-SCHEMEformer-PPAA-44-e8-S24 21.2 4.1 81.88
-SCHEMEformer-PPAA-44-e8-S24 21.2 3.3 81.88
-SCHEMEformer-PPAA-44-e8-S36 55.2 10 83.20
-SCHEMEformer-PPAA-44-e8-S36 55.2 8.0 83.15
-SCHEMEformer-PPAA-12-e8-S12 21.3 3.8 82.00
-SCHEMEformer-PPAA-12-e8-S12 21.3 3.4 81.98
-SCHEMEformer-PPAA-12-e8-S24 40.0 7.3 82.80
-SCHEMEformer-PPAA-12-e8-S24 40.0 6.5 82.76
-SCHEMEformer-PPAA-12-e8-S36 58.8 10.8 84.00
-SCHEMEformer-PPAA-12-e8-S36 58.8 9.6 83.95
-Table 10: Ablation studies on removing CCA branch during inference for the family of SCHEMEformer-PPAA models on ImageNet-1K dataset.
-CCA Group1 Group2 Group3 Group4 Ensemble Model (%) (%) (%) (%) (%) SCHEMEformer-44-e8-S12 29.63 37.59 37.57 35.89 73.95 SCHEMEformer-44-e8-S12 27.89 37.64 30.02 37.67 75.04 Table 11: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module.
-We train a linear classifier on top of the four feature groups extracted from the final MLP mixer of the transformer block of the network.
-The model trained with CCA forms feature clusters that learn diverse and complementary set of features that can obtain 1% higher validation accuracy than the model trained without CCA.
-[…]
-Backbone CCA RetinaNet 1× Mask R-CNN 1× Used #P $AP$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP_{M}$ $AP_{L}$P $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 SCHEMEformer-44-e8-S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8 Table 13: Ablation study of removing CCA for COCO-17 Object Detection and Instance Segmentation.
-Removing CCA at inference does not impact the AP values as they are identical to the model using CCA at inference.
-$AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
-CCA Semantic FPN #Par FLOPs mIoU (%) 15.5 36.4 40.9 15.5 34.3 40.9 24.8 49.8 42.5 24.8 45.7 42.5 Table 14: Ablation study of removing CCA for Semantic Segmentation results on ADE20K dataset.
+The standard transformer MLP was replaced with the SCHEME module to obtain a new family of SCHEMEformer models with improved performance for classification, detection, and segmentation, for fixed parameters and FLOPs with favorable latency.
+SCHEME was shown to be effective for various ViT architectures, always outperforming similar complexity models with smaller MLP expansion ratios.
+Appendix A Appendix
+Table 9: Comparison with state-of-the-art ViT models on the ImageNet-1K dataset.
+The SCHEME module enhances the accuracy of existing ViTs while maintaining or achieving higher throughput.
+* denotes the results reported in the original paper at 224 input resolution.
+Model #Par (M) FLOPs (G) T (img/s) Acc (%)
+ViT-Base [11] 86 17.6 112 79.7
+SCHEME-ViT-12-e8-Base 77.1 15.5 130 (+16%) 79.9 (+0.2%)
+DeiT-Tiny [37] 6 1.3 117 74.5
+SCHEME-DeiT-12-e8-Tiny 7.5 1.6 117 (+0%) 76.0 (+1.5%)
+Poolformer-S12 12.0 1.8 166 77.2
+SCHEME-Poolformer-12-e8-S12 16.7 2.6 171 (+3%) 78.5 (+1.3%)
+Metaformer-S12 12.0 1.8 133 78.9
+SCHEMEformer-PPAA-12-e8-S12 11.8 1.8 133 (+0%) 79.7 (+0.8%)
+CAformer-S12 25 4.2 74 82.9
+SCHEME-CAformer-12-e8-S12 23.9 3.6 80 (+8%) 82.9 (+0.0%)
+CoAtNet-0 25.0 4.2 88 81.6
+SCHEME-CoAtNet-12-e8-0 24.0 4.1 110 (+25%) 81.6 (+0.0%)
+CSWin-Tiny 23.0 4.3 20 82.8
+SCHEME-CSWin-12-e8-Tiny 29.1 5.6 21 (+5%) 83.2 (+0.4%)
+DaViT-Tiny 23.0 4.3 61 82.8
+SCHEME-DaViT-12-e8-Tiny 37.0 6.6 62 (+1%) 83.0 (+0.2%)
+T2T-ViT-14 21.5 6.1 70 81.7
+SCHEME-T2T-ViT-12-e8-14 27.7 8.1 71 (+1%) 82.1 (+0.4%)
+BiFormer-Tiny[61] 13 2.2 57 81.4
+SCHEME-BiFormer-12-e3-Tiny 11.5 1.9 59 (+4%) 81.4 (+0.0%)
+ViT-Large[11] 304 61.6 56 76.5*
+SCHEME-ViT-44-e8-Large 143 28.7 114 (+104%) 81.7 (+5.2%)
+Pyramid-Vision-Llama-S[3] 26 3.3 7 81.6
+SCHEME-Pyramid-Vision-Llama-S 27 3.8 7 (+ 0%) 82.4 (+0.8%)
+A.1 Code Release
+Code is included in the supplementary material. Code and trained models will be released publicly upon acceptance of the paper.
+A.2 Implementation Details
+A.2.1 Hyperparameter Settings
+Table A.2.1 shows the detailed hyperparameter settings of the family of SCHEMEformer models reported in the main paper.
+For reference, SCHEMEformer-44-e8-224 model requires about 3 days to train on 8 NVIDIA TITAN Xp GPUs.
+Table 10: Hyperparameter Settings for the family of SCHEMEformer models trained on ImageNet-1K dataset.
 Model SCHEMEformer-PPAA-44-e8 SCHEMEformer-PPAA-12-e8 SCHEME-CAformer
@@
 Batch size used in the paper 1024
-Peak learning rate used in the paper $4\times 10^{-4}$
 Learning rate decay cosine
@@
 Adam ($\beta_{1}$, $\beta_{2}$) (0.9, 0.999) None
-Table 15: Hyperparameter Settings for the family of SCHEMEformer models trained on ImageNet-1K dataset.
-A.1 Plot of the learned $1-\alpha$ weights in the SCHEME module for SCHEMEformer models
 (a) SCHEMEformer-44-e8-S12
@@
 We demonstrate that these peak weights are not yet converged and training the network for more epochs decays these weights to zero while also improving the accuracy.
-For example, training the SCHEMEformer-PPAA-12-e8-S12 model for 200 additional epochs reduced the weight norm of the vector of $1-\alpha$ weights from 0.18 to 0.05 showing that these weights gradually approach zero as the training progresses while improving the accuracy further by 0.3%.
-The learned weights $1-\alpha$ for the family of SCHEMEformer models is shown in Fig.
-5.
+For example, training the SCHEMEformer-PPAA-12-e8-S12 model for 200 additional epochs reduced the weight norm of the vector of $1-\alpha$ weights from 0.18 to 0.05 showing that these weights gradually approach zero as the training progresses while improving the accuracy further by 0.4%.
+A.3 Ablation Studies
+A.3.1 Plot of the learned 1-$\alpha$ weights in the SCHEME module for SCHEMEformer models
+The learned weights 1-$\alpha$ for the SCHEMEformer model family are shown in Fig.
+6.
 Interestingly, the learned weights coarsely approximate the shape of a gaussian distribution.
 The learned weights reach a peak value in the middle layers of the network and drop to zero for all the other layers.
-The middle of the network typically correspond to the initial few layers of the third stage of the model that contains the maximum number of transformer blocks for all the models shown in Fig.
-5.
+The middle of the network typically corresponds to the initial few layers of the third stage of the model that contains the maximum number of transformer blocks for all the models shown in Fig.
+6.
 We conjecture that the weights for these layers have not fully converged and that feature mixing can still be useful for these layers and so training for more epochs will allow the 1-$\alpha$ weights of these layers to converge to zero.
-To test this hypothesis, we trained the SCHEMEformer-PPAA-12-e8-S12 model for an additional 200 epochs beyond the standard 300 epochs and observed that the peak value of the $1-\alpha$ weights decreased further by 0.11 as compared to the model trained for 300 epochs and the accuracy improved by 0.3%.
+To test this hypothesis, we trained the SCHEMEformer-PPAA-12-e8-S12 model for an additional 200 epochs beyond the standard 300 epochs and observed that the peak value of the 1-$\alpha$ weights decreased further by 0.11 as compared to the model trained for 300 epochs and the accuracy improved by 0.4%.
 The weight norm (of all layers) decreased from 0.18 for 300 epoch model to 0.05 for 500 epoch model.
 This confirms our hypothesis and training SCHEMEformer models for larger epochs can further improve the accuracy.
-A.2 More results of removing CCA at inference
-Table 10 shows the results of the full family of SCHEMEformer models with and without using CCA at inference.
-It is straightforward to observe that removing CCA does not impact the accuracy consistently across different model depths and larger models have slightly higher accuracy drop as compared to smaller models while still being negligible overall ($\leq 0.05\%$).
-Table 13 and 14 show the results of removing CCA at inference for object detection and semantic segmentation models, respectively.
-For both tasks, results are identical to the model using CCA showing that the result also generalizes to downstream tasks.
-A.3 CCA
-In the paper, feature groups across different layers of the model was used to demonstrate the learning of feature clusters by CCA.
+Table 11: Training Overhead of CCA. CCA adds only a small overhead in GPU memory and training time.
+CCA #Par Train FLOPs Val Acc GPU Mem. Train Throughput
+Model (M)$\downarrow$ (G)$\downarrow$ (%) $\uparrow$ (G) $\downarrow$ (iters/s) $\uparrow$
+SCHEMEformer-PPAA-44-e8-S12 11.8 1.77 79.1 8 215
+SCHEMEformer-PPAA-44-e8-S12 11.8 2.16 79.7 9 180
+Table 12: Ablation study of longer training for SCHEMEformer-PPAA-44-e8-S12.
+Model #P (M) FLOPs (G) Throughput (img/s) 300-Acc (%) 500-Acc (%)
+Metaformer-11-e2-S12 (Baseline) 11.8 1.77 133 78.9 79.6
+SCHEMEformer-44-e8-S12 11.8 1.77 133 79.7 80.1
+A.3.2 Training Overhead of CCA
+Table A.3.1 compares the training GPU memory and throughput for SCHEME mixer with and without using CCA.
+CCA improves the accuracy by 0.6% with only a slight increase in the GPU memory (+12.5%) and training time (+16.7%).
+Further, CCA is not needed during inference thereby providing gains for "free" without additional computational cost at inference.
+A.3.3 Training for longer epochs
+Table A.3.1 shows the comparison of training for longer epochs for SCHEMEformer with the baseline Metaformer model.
+We train for an additional 200 epochs from the standard 300 epochs.
+SCHEMEformer-PPAA-44-e8-S12 trained for 300 epochs even outperforms the baseline model trained for 500 epochs.
+On continuing the training from 300 to 500 epochs, SCHEMEformer continues to improve the performance without saturation suggesting that it is beneficial to train with CCA for longer epochs.
+300 to 500 epochs is a much larger increase of training time (67%) than the 16.7% increase in training time required by SCHEME mixer (see Table A.3.1).
+Table 13: Impact of removing CCA branch during inference.
+CCA #Params FLOPs Top-1 Acc
+Model Used  (M) (G) (%)
+SCHEMEformer-PPAA-44-e8-S12 11.83 2.16 79.74
+SCHEMEformer-PPAA-44-e8-S12 11.83 1.77 79.72
+SCHEMEformer-PPAA-12-e8-S24 40.0 7.3 82.80
+SCHEMEformer-PPAA-12-e8-S24 40.0 6.5 82.76
+SCHEMEformer-PPAA-12-e8-S36 58.8 10.8 84.00
+SCHEMEformer-PPAA-12-e8-S36 58.8 9.6 83.95
+SCHEME-CoatNet-44-e8-0 17.80 3.83 80.70
+SCHEME-CoatNet-44-e8-0 17.80 3.42 80.68
+SCHEME-Swin-12-e8-T 36.93 7.00 81.69
+SCHEME-Swin-12-e8-T 36.93 5.89 81.69
+Table 14: Ablation study of removing CCA for COCO-17 Object Detection and Instance Segmentation.
+Removing CCA at inference does not impact the AP values as they are identical to the model using CCA at inference.
+$AP^{b}$ and $AP^{m}$ denote bounding box AP and mask AP, respectively.
+Backbone models denote SCHEMEformer-PPAA-44-e8- variants.
+Backbone CCA RetinaNet 1× Mask R-CNN 1×
+Used #P $AP$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP_{M}$ $AP_{L}$P $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3
+S12 21 38.3 58.0 40.4 21.0 41.4 52.3 31 39.8 61.9 42.9 24.1 53.0 42.3
+S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8
+S24 31 38.8 58.7 41.2 22.5 41.5 53.5 41 40.9 62.5 44.6 24.6 55.6 43.8
+Table 15: Ablation study of removing CCA for Semantic Segmentation results on ADE20K validation dataset.
+CCA Semantic FPN
+#Par FLOPs mIoU (%)
+15.5 36.4 40.9
+15.5 34.3 40.9
+24.8 49.8 42.5
+24.8 45.7 42.5
+A.3.4 Impact of removing CCA at inference
+Table A.3.3 shows the impact of removing CCA at inference, for various backbones.
+While the number of FLOPS decreases, the top-1 accuracy changes very little ($\approx 0.02$ difference).
+Hence, there is no advantage in using CCA at inference.
+This is unlike training, where the use of CCA makes a non-negligible difference, as shown in Table 6 of the main paper.
+Table A.3.3 and A.3.3 show the results of removing CCA at inference for object detection and semantic segmentation models, respectively.
+For both tasks, the results are identical to the model using CCA showing that the CCA also generalizes to downstream tasks.
+Table 16: Ablation study on the effect of larger expansion ratios in BD-MLP block of SCHEME on ImageNet-1K validation dataset.
+#Par FLOPs Top-1
+Model (M) (G) Acc (%)
+Metaformer-PPAA-11-e1-S12 9.6 1.37 76.0
+SCHEMEformer-PPAA-22-e2-S12 9.6 1.37 77.0
+SCHEMEformer-PPAA-44-e4-S12 9.6 1.37 77.4
+SCHEMEformer-PPAA-66-e6-S12 10.0 1.45 77.9
+Metaformer-PPAA-11-e2-S12 11.8 1.77 78.9
+SCHEMEformer-PPAA-22-e4-S12 11.8 1.77 79.4
+SCHEMEformer-PPAA-44-e8-S12 11.8 1.77 79.7
+SCHEMEformer-PPAA-33-e6-S12 12.5 1.87 79.8
+SCHEMEformer-PPAA-22-e6-S12 14.1 2.16 80.4
+Metaformer-PPAA-11-e4-S12 16.5 2.56 81.0
+SCHEMEformer-PPAA-22-e8-S12 16.5 2.56 81.1
+SCHEMEformer-PPAA-44-e16-S12 16.5 2.56 81.2
+Table 17: Ablation study on the formation of feature clusters in the BD-MLP branch of the SCHEME module.
+We train a linear classifier on top of the four feature groups extracted from the final MLP mixer of the transformer block of the network.
+The model trained with CCA forms feature clusters that learn diverse and complementary set of features that can obtain 1% higher validation accuracy than the model trained without CCA.
+CCA Group1 Group2 Group3 Group4 Ensemble
+Model (%) (%) (%) (%) (%)
+SCHEMEformer-44-e8-S12 29.63 37.59 37.57 35.89 73.95
+SCHEMEformer-44-e8-S12 27.89 37.64 30.02 37.67 75.04
+Figure 7: GRAD-CAM [33] visualization for a few validation samples on ImageNet-1K dataset for SCHEMEformer-PPAA-44-e8-S12 model and comparison with other competing methods.
+A.3.5 CCA
+In the main paper, feature groups across different layers of the model was used to demonstrate the learning of feature clusters by CCA.
 Here, the feature from the final layer of the model is only considered.
-Table 11 shows the top-1 accuracy per feature group and model.
+Table A.3.4 shows the top-1 accuracy per feature group and model.
 The average of the outputs from the four group classifiers is reported in the final column of the table.
-The effect is more pronounced when using a single layer feature with more than 1% accuracy difference between the model with and without CCA.
+The effect is more pronounced when using a single layer feature with +1.09% accuracy difference between the model with and without CCA.
 This further reinforces that by introducing inter-group communication, CCA enables the groups to learn more diverse sets of features, that complement each other.
-A.4 Larger expansion ratios
-Table 12 shows additional results of using larger expansion ratios with SCHEME for the same parameters and FLOPs using the Metaformer-PPAA-S12 baseline model.
-We observe that SCHEMEformer-PPAA consistently outperforms the baseline for larger expansions ratios and saturates as the model size and FLOPs increases.
-Appendix B Implementation Details
-B.1 Hyperparameter Settings
-Table 15 shows the detailed hyperparameter settings of the family of SCHEMEformer models reported in the main paper.
-Appendix C Qualitative Analysis: Grad-CAM [38] Visualization
+A.3.6 Larger expansion ratios
+Table A.3.4 shows additional results of using larger expansion ratios with SCHEME mixer (illustrated in Fig.
+3 of the main paper) for the same number of parameters and FLOPs using the Metaformer-PPAA-S12 baseline model.
+We observe that SCHEMEformer-PPAA consistently outperforms the baseline for larger expansions ratios with larger gains at lower FLOPs.
+The performance saturates as the model size and FLOPs increases.
+Figure 8: Visualization of last layer attention maps of the SCHEMEformer-PPAA-44-e8-S12 model with and without CCA on Imagenet validation dataset images.
+The ellipses identify the region of the object in the target class.
+Since the SCHEMEformer model does not have class tokens, the attention maps are not necessarily interpretable.
+The spikes in the attention maps are also common in attention visualizations and can be avoided by using registers [7].
+A.4 Qualitative Analysis
+A.4.1 Grad-CAM [33] Visualization
 Fig.
-6 shows the results of class activation maps for SCHEMEformer-PPAA-44-e8-S12 model for few examples from the validation set of ImageNet-1K dataset.
+7 shows the results of class activation maps for SCHEMEformer-PPAA-44-e8-S12 model for a few examples from the validation set of ImageNet-1K dataset.
 The stronger heatmap responses around the salient features of an object (e.g., body of a bird, cat) shows that the model ignores the background and attends to more discriminative spatial regions.
 Fig.
-6 also shows the comparison with a few existing methods that demonstrates that SCHEMEformer-PPAA-44-e8-S12 attends to the complete object class and is better than the competing methods.
-(a) Input Images
-(b) ResNet-50 [21]
-(c) ResMLP [42]
-(d) DeiT-Small[43]
-(e) Poolformer-S12[57]
-(f) SCHEMEformer-PPAA-44-e8-S12
-Figure 6: GRAD-CAM [38] visualization for a few validation samples on ImageNet-1K dataset for SCHEMEformer-PPAA-44-e8-S12 model.
+7 also shows the qualitative comparison with a few existing methods such as ResNet-50, DeiT-S, Poolformer etc.
+of similar complexity.
+It demonstrates that SCHEMEformer-PPAA-44-e8-S12 attends to the complete object class and less spurious features showing that it is better than the competing methods.
+A.4.2 Attention Visualization
+Figure 8 shows the results of attention maps with and without CCA on Imagenet dataset.
+We find that there are differences in attention between using or not using CCA.
+CCA increases the tendency of the model to attend to the regions where the class is present, which is denoted by green ellipses.
+(a) Accuracy-FLOPs
+(b) Accuracy-Parameters
+(c) Accuracy-Throughput
+Figure 9: [Zoomed version] Comparison of the proposed SCHEMEformer family, derived from the Metaformer-PPAA-S12 model [55] with higher expansion ratios in the MLP blocks, and many SOTA transformers from the literature.
+The SCHEMEFormer family establishes a new Pareto frontier (optimal trade-off) for a) accuracy vs.
+FLOPs, b) accuracy vs model size, and c) accuracy vs, throughput.
+SCHEMEformer models are particularly effective for the design of fast transformers (throughput between 75 and 150 images/s) with small model size.
+Figure 10: [Zoomed version] SCHEME tradeoffs.
+Top: Accuracy vs FLOPs of various SCHEME models with different MLP configurations.
+Bottom: SCHEME mixer improves accuracy for fixed throughput or vice-versa for various popular ViT architectures.
+Figure 11: [Zoomed version] Impact of CCA (SCHEMEformer-44-e8-S12).
+Top: Evolution of weight $1-\alpha$ across model layers.
+Bottom: Class separability of output features $\bf y$ (over 50 random classes of ImageNet-1K validation set) for model trained with and without CCA.